### v1

# Guess your neighbor’s input: Quantum advantage in Feige’s game

Abstract.
In this article, we study a nonlocal game with two questions and three answers per player, which was first considered by Feige in 1991, and show that there is quantum advantage in this game. We prove that the game is a robust self-test for the 33-dimensional maximally entangled state. Furthermore, we show that the game can be seen as the ”or” of two games that each do not have quantum advantage. Lastly, we investigate the behavior of the game with respect to parallel repetition in the classical, quantum and non-signalling case and obtain perfect parallel repetition of the non-signalling value if Feige’s game is repeated an even amount of times.
1. Introduction
A nonlocal game CHTW (04) consists of two (or more) cooperative players, often called Alice and Bob, that interact with a referee. In the game, the referee sends a question to each player, for which each player gives an answer. Based on the tuple of questions and answers, the referee decides if the players win or lose. Communication is not permitted between Alice and Bob, hence each player has no information about the questions given to the other players, nor do they know the answers provided to the referee by the other players. Nevertheless, the description of the game is known to the players ahead of time, allowing them to agree on a strategy beforehand and maximize their probability of winning the game. The classical value $\omega_{c}(G)$ of a nonlocal game $G$ is the maximum winning probability of classical players, and the quantum value $\omega_{q}(G)$ denotes the maximum winning probability of quantum players sharing a (finite) amount of quantum resources (such as entangled quantum states, like EPR pairs). Furthermore, the non-signalling value $\omega_{ns}(G)$ is the maximum winning probability if one considers probability distributions such that the marginal distribution of either player’s answers must be independent of the other player’s question, without requiring the probability distribution to be physically realizable.
Feige’s game. In this work, we revisit a game first studied by Feige Fei (91), where it was considered as counterexample in parallel repetition since its value does not decrease when played twice in parallel. The game is defined as follows. Alice and Bob both receive a bit as question (with uniform probability), and they are each allowed to choose an answer from the set $\{0,1,\perp\}$. To win the game, exactly one of the players has to answer with $\perp$, while the other player’s answer has to agree with the question the player who answered $\perp$ received. More formally, the game $G_{F}$ Fei (91) is defined via
$X=Y=\{0,1\},\quad A=B=\{0,1,\perp\},\quad\pi(x,y)=\tfrac{1}{4}\quad\text{for all }(x,y)$
and has the predicate
$V(a,b,x,y)=\begin{cases}1\quad(a,b)=(\perp,x)\text{ or }(a,b)=(y,\perp),\\ 0\quad\text{otherwise.}\end{cases}$
It has been conjectured in the literature that $\omega_{q}(G_{F})=\omega_{c}(G_{F})$ (Yue (16)), but so far no attempts for computing the quantum value of the game were made. Since there is no quantum advantage for other variants of ”guess your neighbors input” games (ABB+ (10), AAAB (16)), it sounds plausible that there is also no quantum advantage in Feige’s game. We show that, surprisingly, there exists a quantum strategy for Feige’s game that exceeds the classical value and we compute the quantum value.
Theorem.
Feige’s game $G_{F}$ has quantum advantage with $\omega_{q}(G_{F})=\frac{9}{16}>\frac{1}{2}=\omega_{c}(G_{F})$.
Interestingly, the players use operators acting on $\mathbb{C}^{3}$ in the optimal quantum strategy and they share the maximally entangled state $\mathinner{|\psi_{3}\rangle}=\sum_{i=0}^{2}\mathinner{|i\rangle}\otimes\mathinner{|i\rangle}$. The strategy that achieves a winning probability of $\frac{9}{16}$ is constructed from the operators
$\tilde{Z}=Z\oplus 1\quad\text{and}\quad\tilde{X}=X\oplus 1\quad\text{on}\quad\mathbb{C}^{3}=\mathbb{C}^{2}\oplus\mathbb{C},$
where $X$ and $Z$ denote the usual Pauli matrices. The projective measurements are given by
$P^{x}_{a}=\tilde{X}{\tilde{Z}}^{x}\tilde{X}\mathinner{|\tilde{a}\rangle}\mathinner{\langle\tilde{a}|}\tilde{X}{\tilde{Z}}^{x}\tilde{X},\quad a\in\{0,1,\bot\}\quad\text{ and}$ $Q^{y}_{b}=\tilde{X}P^{y}_{b}\tilde{X},\quad b\in\{0,1,\bot\},$
for a carefully chosen basis $\{\mathinner{|\tilde{a}\rangle}\}$ of $\mathbb{C}^{3}$. To obtain this basis, we crucially use that $\tilde{Z}$ and $\tilde{X}$ are two non-commuting observables such that $\tilde{Z}$ and $\tilde{X}\tilde{Z}\tilde{X}$ commute.
An approach to upper-bounding the quantum value of a nonlocal game is to associate a formal game polynomial $p$ to a game $G$, whose variables are identified with abstract projective measurements $P^{x}_{a}$ and $Q^{y}_{b}$ such that $\mathinner{\langle\psi|}p(P^{x}_{a},Q^{y}_{b})\mathinner{|\psi\rangle}$ is precisely the winning probability of a quantum (commuting) strategy for $G$. To prove that the quantum value of $G_{F}$ is at most $\frac{9}{16}$, we find a sums of square decomposition of the game polynomial of Feige’s game by using the computer. More precisely, we obtain a vector of formal variables $V=V(P^{x}_{a},Q^{y}_{b})$ and a positive complex matrix $Y$ such that $V^{*}YV=\frac{9}{16}I-p(P^{x}_{a},Q^{y}_{b})$, which implies $\omega_{q}(G_{F})\leq\frac{9}{16}$.
Self-testing. A nonlocal game is a self-test if its optimal quantum strategy is unique up to local isometries, as introduced in MY (04). Self-testing is a useful property of a nonlocal game as it can be used for certifying quantum devices. We have the following theorem for Feige’s game.
Theorem.
The game $G_{F}$ is a robust self-test for the maximally entangled state $\mathinner{|\psi_{3}\rangle}=\sum_{i=0}^{2}\mathinner{|i\rangle}\otimes\mathinner{|i\rangle}$.
While self-tests for all maximally entangled states are known (MPS (24); Fu (22)), we believe that this is the smallest game that self-tests $\mathinner{|\psi_{3}\rangle}$ with just two questions and three answers per player.
In fact, we obtain the result by showing that a certain interesting algebra has a unique irreducible representation. We deduce the algebra from the sums of square decomposition we have seen before. Since it holds $V^{*}YV=0$ in an optimal strategy, there are relations that the operators in the strategy have to fulfill. In our case we get the relations for Alice’s operators
$P^{x}_{a}P^{1-x}_{a^{\prime}}P^{x}_{a}\mathinner{|\psi\rangle}$ $=\nu_{aa^{\prime}}P^{x}_{a}\mathinner{|\psi\rangle}\text{ for all }x\in\{0,1\},a,a^{\prime}\in\{0,1,\bot\},\text{ where }\nu=\begin{bmatrix}\frac{9}{16}&\frac{1}{16}&\frac{3}{8}\\ \frac{1}{16}&\frac{9}{16}&\frac{3}{8}\\ \frac{3}{8}&\frac{3}{8}&\frac{1}{4}\end{bmatrix}$.
The same relations hold for Bob’s operators. Note that the relations look similar to the relations obtained from mutually unbiased measurements TFR+ (21), where for the latter it holds $\nu=\frac{1}{d}I$. In case of $P^{x}_{a}=\mathinner{|\phi^{x}_{a}\rangle}\mathinner{\langle\phi_{a}^{x}|}$ being rank 11 projections, the transition matrix from the basis $\{\mathinner{|\phi^{x}_{a}\rangle}\}$ to the basis $\{\mathinner{|\phi^{1-x}_{a}\rangle}\}$ is given by a unitary matrix $H$ with $|h_{aa^{\prime}}|=\sqrt{\nu_{aa^{\prime}}}$ (which is a Hadamard matrix in the case of $\nu=\frac{1}{d}I$). Similar to some Hadamard matrices being unique up to equivalence Haa (97), we show that the same is true for unitaries with $|h_{aa^{\prime}}|=\sqrt{\nu_{aa^{\prime}}}$ for the specific choice of $\nu$ from above. More generally, we give simple criterion for when such a unitary is unique up to a certain kind of equivalence. Using this, and the operator-algebraic framework for self-testing developed in Zha (24), we obtain that Feige’s game is a self-test. Another ingredient we need in our proof is that in an optimal strategy, the actions of Bob’s projective measurements on the state are determined by actions of (self-adjoint polynomials in) Alice’s projective measurements.
Feige’s game as ”or”-game. Feige’s game has another interesting property, it can be seen as the ”or” of two games (MS (23)). For two nonlocal games $G_{1}$ and $G_{2}$, the $(G_{1}\lor G_{2})$-game is defined as follows. The referee sends Alice and Bob a pair of questions $(x_{1},x_{2})$ and $(y_{1},y_{2})$, respectively, where $x_{i},y_{i}$ are questions in $G_{i}$, $i\in\{1,2\}$. Each of them chooses one of the questions they received and responds with an answer from the corresponding game. To win the game, two conditions have to be fulfilled:
  • (1)
    Alice and Bob have to give answers from the same game,
  • (2)
    their answers have to win the corresponding game.
Since in Feige’s game, one of the players has to guess the input of the other player, it can be realized as the ”or”-game of the following two games: In the first game, Alice always guesses the input of Bob, while he outputs $\perp$ all the time and in the second game the roles are reversed. It is intriguing to note that the games for which a fixed player guesses the input of the other player do not have quantum advantage. We therefore have an example of an ”or”-game whose quantum value is strictly higher than the maximum of the quantum values of the games it is constructed from.
Theorem.
There are nonlocal games $G_{1}$ and $G_{2}$ such that $G_{F}=G_{1}\lor G_{2}$ and
$\omega_{q}(G_{F})=\frac{9}{16}>\frac{1}{2}=\mathrm{max}\{\omega_{q}(G_{1}),\omega_{q}(G_{2})\}.$
This theorem is in contrast to the case of perfect quantum strategies (MS (23)). Here it is known that if the players give answers for one of the games in a perfect strategy of the ”or”-game, then this game has to have a perfect quantum strategy as well.
Parallel repetition. The original motivation of Feige was to give a counterexample in parallel repetition as its value does not decrease when played twice in parallel. In the $n$-fold parallel repetition $G^{\times n}$ of a game $G$, the players receive tuples of questions $(x_{1},\dots,x_{n})$ and $(y_{1},\dots,y_{n})$ from the original game and they answer with tuples $(a_{1},\dots,a_{n})$ and $(b_{1},\dots,b_{n})$. Alice and Bob win in $G^{\times n}$ if the questions and answers $x_{i},y_{i},a_{i},b_{i}$ win the game $G$ for each $i\in\{1,\dots,n\}$. We study the behaviour of the game with respect to parallel repetition for the classical, quantum and non-signalling value. The results are summarized in Table 1.
$n=1$ $n=2$ $n=3$ $n$ even
$\omega_{c}$ $\tfrac{1}{2}$ $\tfrac{1}{2}$ $\mathbf{\tfrac{5}{16}}$ $\tfrac{1}{2^{n/2}}$
$\omega_{q}$ $\mathbf{\tfrac{9}{16}}$ $\mathbf{\tfrac{1}{2}}$ ? $\mathbf{\tfrac{1}{2^{n/2}}}$
$\omega_{ns}$ $\mathbf{\tfrac{2}{3}}$ $\mathbf{\tfrac{1}{2}}$ $\mathbf{\tfrac{1}{{3}}}$ $\mathbf{\tfrac{1}{2^{n/2}}}$
Table 1. Summary of values of the $n$-fold parallel repetition of Feige’s game, where we highlight the values that were previously not known.
In the table, the classical, quantum and non-signalling values agree in the case of an even number of repetitions, which follows from the following theorem.
Theorem.
It holds $\omega_{ns}(G_{F}^{\times n})=\frac{1}{2^{n/2}}$ for $n$ even.
Note that this is not the case for odd numbers. From the previous discussion, we know that the classical value does not agree with the quantum value for the original game, and we show that the non-signalling value is even higher. Furthermore, we get that at least the classical and non-signalling value are different for $n=3$, despite this not being the case for $n=2$. We do not know if there is quantum advantage for $n=3$.
Question.
Does the quantum value exceed the classical value for the 33-fold parallel repetition of Feige’s game?
We are not aware of a game with quantum advantage, for which the classical and quantum value agree for the 22-fold parallel repetition, but then they separate again for $n=3$. As we can see in Table 1 this happens for the classical and the non-signalling value of Feige’s game. Note that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition performs worse than the optimal classical strategy for $n=3$, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$. Therefore, there is no natural candidate of a quantum strategy that yields quantum advantage for the 33-fold parallel repetition. The $\frac{5}{16}$ optimal classical strategy is also interesting: It performs worse on the first two rounds of the game in comparison to the optimal strategy for the 22-fold parallel repetition, but is constructed such that if the players win the first two rounds, they win the third round automatically.
2. Preliminaries
Let $A,B,X,Y$ be finite sets, $\pi:X\times Y\to[0,1]$ a probability distribution and $V:A\times B\times X\times Y\to\{0,1\}$ a function.
Definition 2.1.
A (two-player) nonlocal game is a tuple $G=(X,Y,A,B,\pi,V)$ describing a scenario consisting of non-communicating players, Alice and Bob, interacting with a referee. In the game, the referee samples a pair of questions $(x,y)\in X\times Y$ according to $\pi$, sending question $x$ to Alice and $y$ to Bob. Then, Alice (respectively Bob) returns answer $a$ (respectively $b$) to the referee. The players win if $V(a,b|x,y)=1$, otherwise they lose.
The players are not allowed to communicate during the game, but they can agree on a strategy beforehand. For each strategy, we obtain probabilities $p(a,b|x,y)$ of Alice and Bob answering $a,b$ given $x,y$. The collection of numbers $\{p(a,b|x,y)\}$ is called a correlation. The winning probability for the game, using this strategy, is given by
$\omega(G,S)=\omega(G,p)=\sum_{a,b,x,y}\pi(x,y)V(a,b|x,y)p(a,b|x,y)$
2.1. Strategies for nonlocal games
There are different types of strategies, which we now describe. We start with classical strategies.
Definition $2.$
A classical strategy for a nonlocal game $G$ consists of
  1. (1)
    a probability distribution $\gamma\colon\Omega\to[0,1]$ on a (without loss of generality) finite probability space $\Omega$, along with
  2. (2)
    probability distributions $\{p_{\omega}(a|x):x\in X,\omega\in\Omega\}$ with outcomes in AA and $\{q_{\omega}(b|y):y\in Y,\omega\in\Omega\}$ with outcomes in $B$.
In this case, we have
$p(a,b|x,y)=\sum_{\omega\in\Omega}\gamma(\omega)\,p_{\omega}(a|x)\,q_{\omega}(b|y).$
A special case of classical strategies are deterministic strategies, in which the answers of the players are determined by the questions. In this case, the strategies are described by functions $f:X\to A$ and $g:Y\to B$, and Alice and Bob answer with $f(x)$ and $g(y)$ given $x$ and $y$, respectively.
In a quantum strategy, the players are allowed to share a quantum state and perform local measurements. A (quantum) state $\mathinner{|\psi\rangle}$ is a unit vector in a Hilbert space $\mathcal{H}$. The measurements are described by positive operator valued measures (POVMs), which consist of a family of positive operators $\{M_{i}\in B(\mathcal{H})\,|\,i\in[m]\}$ such that $\sum_{i=1}^{m}M_{i}=1_{B(\mathcal{H})}$. If all positive operators are projections ($M_{i}=M_{i}^{*}=M_{i}^{2}$), then we call $\{M_{i}\in B(\mathcal{H})\,|\,i\in[m]\}$ with $\sum_{i=1}^{m}M_{i}=1_{B(\mathcal{H})}$ a projective measurement (PVM).
Definition 2.3.
A quantum strategy for a nonlocal game $G$ consists of
  1. (1)
    a (without loss of generality) pure quantum state $\mathinner{|\psi\rangle}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, along with
  2. (2)
    POVMs $\{\{P_{a}^{x}:a\in A\}:x\in X\}$ acting on $\mathcal{H}_{A}$ and POVMs $\{\{Q_{b}^{y}:b\in B\}:y\in Y\}$ acting on $\mathcal{H}_{B}$.
Then
$p(a,b|x,y)=\mathinner{\langle\psi|}P_{a}^{x}\otimes Q_{b}^{y}\mathinner{|\psi\rangle}.$
Now, we will define non-signalling strategies. These strategies do not have to be physical, but the resulting probability distribution ensures that Alice does not use Bob’s input for her answer and vice versa.
Definition 2.4.
A non-signalling strategy consists of a correlation $\{p(a,b|x,y)\}$ such that
$\sum_{a\in A}p(a,b|x,y)$ $=\sum_{a\in A}p(a,b|\tilde{x},y):=p(b|y),$
$\sum_{b\in B}p(a,b|x,y)$ $=\sum_{b\in B}p(a,b|x,\tilde{y}):=p(a|b)$
for all $x,\tilde{x}\in X$ and $y,\tilde{y}\in Y$.
It is natural to ask how well players can perform in a nonlocal game for a specific class of strategies. Therefore, one defines the value of a game as the highest winning probability for a given class of strategies.
Definition 2.5.
For a nonlocal game $G$ and we define
$\omega_{c}(G)$ $=\sup_{S\text{ classical strategy}}\omega(G,S),$
$\omega_{q}(G)$ $=\sup_{S\text{ quantum strategy}}\omega(G,S),$
$\omega_{ns}(G)$ $=\sup_{S\text{ non-signalling strategy}}\omega(G,S).$
We call $\omega_{c}(G)$ the classical value, $\omega_{q}(G)$ the quantum value and $\omega_{ns}(G)$ the non-signalling value of the game.
From the definition, it is immediate that
$\omega_{c}(G)\leq\omega_{q}(G)\leq\omega_{ns}(G).$
Note that there exist games for which all these values are different, for example the CHSH game CHSH (69).
The proof of our main theorem follows an operator-algebraic framework for self-testing developed inZha (24). For finite sets $X$ and AA, we use $\mathscr{A}_{PVM}^{X,A}$ to denote the universal $C^{*}$-algebra generated by $\{p^{x}_{a}:a\in A,x\in X\}$, subject to the relations
  • $(p^{x}_{a})^{2}=(p^{x}_{a})^{*}=p^{x}_{a}\quad\forall a\in A,x\in X$
  • $\sum_{a\in A}p^{x}_{a}=1\quad\forall x\in X$
Given a nonlocal game $G=(X,Y,A,B,\mu,V)$, for any strategy
$S=(\mathinner{|\psi\rangle}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B},\{P^{x}_{a}\},\{Q^{y}_{b}\}),$ (1)
by the universal property, there are unique *-representations
$\pi_{A}:\mathscr{A}_{PVM}^{X,A}\rightarrow\mathcal{B}(\mathcal{H}_{A})\text{ and }\pi_{B}:\mathscr{A}_{PVM}^{Y,B}\rightarrow\mathcal{B}(\mathcal{H}_{B})$
such that
$\pi_{A}(p^{x}_{a})=P^{x}_{a}\text{ and }\pi_{B}(q^{y}_{b})=Q^{y}_{b}\quad\forall a\in A,b\in B,x\in X,y\in Y.$ (2)
In the following we reserve the symbols $p^{x}_{a}$ and $q^{y}_{b}$ for the generators for $\mathscr{A}_{PVM}^{X,A}$ and $\mathscr{A}_{PVM}^{Y,B}$ respectively, whereas $P^{x}_{a}$ and $Q^{y}_{b}$ will denote representations given by operators on Hilbert spaces as in (2). We call $\pi_{A},\pi_{B}$ the representations associated with $S$ as in (1), and sometimes write
$S=(\mathinner{|\psi\rangle}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B},\pi_{A},\pi_{B})$ (3)
for a strategy.
Given a nonlocal game $G=(X,Y,A,B,\mu,V)$, we define the game polynomial as
$\Phi_{G}:=\sum_{a,b,x,y}\mu(x,y)V(a,b|x,y)p^{x}_{a}\otimes q^{y}_{b}$
which is an element of $\mathscr{A}_{PVM}^{X,A}\otimes\mathscr{A}_{PVM}^{Y,B}$ that encodes all the information of $G$. The value at the strategy (3) is then given by
$\omega(G,S)=\mathinner{\langle\psi|}\pi_{A}\otimes\pi_{B}(\Phi_{G})\mathinner{|\psi\rangle}.$
3. Feige’s game
In this section, we will discuss the classical, quantum and non-signalling value of a game that appeared in Fei (91). The game was considered as counterexample in parallel repetition, as its classical value does not decrease when played twice in parallel. As main result of this section, we will show that its quantum value exceeds the classical value.
We recall again the definition of the game from the introduction. Alice and Bob both receive a bit as question (with uniform probability), and they are each allowed to choose an answer from the set $\{0,1,\perp\}$. To win the game, exactly one of the players has to answer with $\perp$, while the second player has to guess the input of the first player, i.e. the answer is a bit that is equal to the question of first player. More formally, the game $G_{F}$ Fei (91) is defined via
$X=Y=\{0,1\},\quad A=B=\{0,1,\perp\},\quad\pi(x,y)=\tfrac{1}{4}\quad\forall(x,y)$
and has the predicate
$V(a,b,x,y)=\begin{cases}1\quad(a,b)=(\perp,x)\text{ or }(a,b)=(y,\perp),\\ 0\quad\text{otherwise.}\end{cases}$
We will now argue that the classical value of this game is $\frac{1}{2}$. It is well-known that the classical value of a game is attained by a deterministic strategy. It is easy to see that the strategy in which Alice always answers $\perp$ and Bob always answers 0 has winning probability $\frac{1}{2}$, since Alice receives the question 0 halve of the time.
Note that in a deterministic strategy, winning or losing just depends on the input of the players. We have the following observations.
  • (i)
    Once a player answers a bit given a question (e.g. Alice answers 0 when receiving 0), the game will always be lost if the other player receives a question different from this bit (the referee sends 11 to Bob), which happens halve of the time.
  • (ii)
    In a deterministic strategy, the case that exactly one player answers $\perp$ and the other player answers with a bit is only possible for zero, two or all four different input pairs $\{(0,0),(0,1),(1,0),(1,1)\}$. If this happens for zero or two input pairs, the winning probability is upper bounded by $\frac{1}{2}$. If it is true for all input pairs, we get by (i) that the winning probability of the strategy is $\frac{1}{2}$.
Summarizing, we obtain $\omega_{c}(G_{F})=\frac{1}{2}$. In the following, we will show that there exists a quantum strategy that allows the players to win the game with a higher probability. We will see that the winning probability of the quantum strategy is $\frac{9}{16}$. We show that this is an optimal quantum strategy for Feige’s game in Appendix A and obtain a self-testing statement in Section 4.
3.1. Construction of a quantum strategy exceeding the classical value
The game polynomial for $G_{F}$ is given by
$\beta(p^{x}_{a},q^{y}_{b})=\tfrac{1}{4}\sum_{x,y\in\{0,1\}}(p^{x}_{\perp}\otimes q^{y}_{x}+p^{x}_{y}\otimes q^{y}_{\perp}).$
Consider the operators
$\tilde{Z}=Z\oplus 1\quad\text{and}\quad\tilde{X}=X\oplus 1\quad\text{on}\quad\mathbb{C}^{3}=\mathbb{C}^{2}\oplus\mathbb{C}$
(i.e., $\mathbb{C}^{2}$ is the defining and $\mathbb{C}$ the trivial representation of the Pauli group). Note that $\tilde{Z}$ and $\tilde{X}$ are two non-commuting observables such that $\tilde{Z}$ and $\tilde{X}\tilde{Z}\tilde{X}$ commute (and every such pair of projections essentially looks like this, up to multiplicities). Choose $\mathinner{|\bot\rangle}\in\operatorname{Fix}(\tilde{Z}){\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\ =\mathinner{|1\rangle}^{\perp}}$, for now arbitrarily, and extend it by $\mathinner{|\tilde{0}\rangle},\mathinner{|\tilde{1}\rangle}$ to a basis of $\mathbb{C}^{3}$. Alice’s and Bob’s strategy is then given by,
$\tilde{S}_{F}=(\mathinner{|\psi\rangle}\in\mathbb{C}^{3}\otimes\mathbb{C}^{3},\{P^{x}_{a}\},\{Q^{y}_{b}\})$
where $\mathinner{|\psi\rangle}=\frac{1}{\sqrt{3}}\left(\mathinner{|00\rangle}+\mathinner{|11\rangle}+\mathinner{|22\rangle}\right)$ and
$P^{x}_{a}$ $=\tilde{X}{\tilde{Z}}^{x}\tilde{X}\mathinner{|\tilde{a}\rangle}\mathinner{\langle\tilde{a}|}\tilde{X}{\tilde{Z}}^{x}\tilde{X}$ $(a\in\{0,1\})$
$P^{x}_{\bot}$ $=\tilde{X}{\tilde{Z}}^{x}\tilde{X}\mathinner{|\bot\rangle}\mathinner{\langle\bot|}\tilde{X}{\tilde{Z}}^{x}\tilde{X}$
$Q^{y}_{b}$ $=\tilde{X}P^{y}_{b}\tilde{X}$ $(b\in\{0,1,\bot\})$
The winning probability is given by
$\omega(G_{F},\tilde{S}_{F})$ $=\frac{1}{4}\sum_{x,y\in\{0,1\}}\frac{1}{3}\operatorname{Tr}[P^{x}_{y}Q^{y}_{\bot}+P^{x}_{\bot}Q^{y}_{x}]=\frac{1}{6}\sum_{x,y\in\{0,1\}}\operatorname{Tr}[P^{x}_{y}Q^{y}_{\bot}]$
$=\frac{1}{6}\sum_{x,y\in\{0,1\}}\operatorname{Tr}[\tilde{X}{\tilde{Z}}^{x}\tilde{X}\mathinner{|\tilde{y}\rangle}\mathinner{\langle\tilde{y}|}\tilde{X}{\tilde{Z}}^{x}\tilde{X}{\tilde{Z}}^{y}\tilde{X}\mathinner{|\bot\rangle}\mathinner{\langle\bot|}\tilde{X}{\tilde{Z}}^{y}]$
$=\frac{1}{3}\sum_{y\in\{0,1\}}\mathinner{\langle\bot|}\tilde{X}{\tilde{Z}}^{y}\mathinner{|\tilde{y}\rangle}\mathinner{\langle\tilde{y}|}{\tilde{Z}}^{y}\tilde{X}\mathinner{|\bot\rangle}$
[…]
$\mathinner{|\bot\rangle}=\sqrt{p}\underbrace{\mathinner{|\text{nontriv}\rangle}}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}=\mathinner{|0\rangle}}+\sqrt{1-p}\underbrace{\mathinner{|\text{triv}\rangle}}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}=\mathinner{|2\rangle}},$
for some $p\in[0,1]$, where we have $\tilde{X}\mathinner{|\text{triv}\rangle}=\tilde{Z}\mathinner{|\text{triv}\rangle}=\mathinner{|\text{triv}\rangle}$, while $\tilde{Z}\mathinner{|\text{nontriv}\rangle}=-\tilde{X}\tilde{Z}\tilde{X}\mathinner{|\text{nontriv}\rangle}=\mathinner{|\text{nontriv}\rangle}$. Then the above winning probability can be computed as follows
$\omega(G_{F},\tilde{S}_{F})$ $=\frac{p}{3}\underbrace{\sum_{y\in\{0,1\}}\mathinner{\langle\text{nontriv}|}\tilde{X}{\tilde{Z}}^{y}\mathinner{|\tilde{y}\rangle}\mathinner{\langle\tilde{y}|}{\tilde{Z}}^{y}\tilde{X}\mathinner{|\text{nontriv}\rangle}}_{=1}$
$\quad+\frac{2\sqrt{p(1-p)}}{3}\sum_{y\in\{0,1\}}\mathinner{\langle\text{triv}|}\tilde{X}{\tilde{Z}}^{y}\mathinner{|\tilde{y}\rangle}\mathinner{\langle\tilde{y}|}{\tilde{Z}}^{y}\tilde{X}\mathinner{|\text{nontriv}\rangle}$
$\quad+\frac{1-p}{3}\underbrace{\sum_{y\in\{0,1\}}\mathinner{\langle\text{triv}|}\tilde{X}{\tilde{Z}}^{y}\mathinner{|\tilde{y}\rangle}\mathinner{\langle\tilde{y}|}{\tilde{Z}}^{y}\tilde{X}\mathinner{|\text{triv}\rangle}}_{=1-(1-p)=p}$
[…]
where we used that
[…]
Note that
$R=\sum_{y\in\{0,1\}}(-1)^{y}\mathinner{|\tilde{y}\rangle}\mathinner{\langle\tilde{y}|}$
[…]
$\begin{split}\mathinner{\langle\text{triv}|}R\tilde{X}\mathinner{|\text{nontriv}\rangle}=&\lVert({\mathds{1}}-|\bot\rangle\langle\bot|)\mathinner{|\text{triv}\rangle}\rVert\\ =&\lVert p\mathinner{|\text{triv}\rangle}+\sqrt{p(1-p)}\mathinner{|\text{nontriv}\rangle}\rVert\\ =&\sqrt{p}.\end{split}$
In this case, the winning probability is
[…]
This is maximized for $p=\frac{3}{4}$, with corresponding success probability
$\omega(G_{F},\tilde{S}_{F})=\frac{9}{16}.$
Concretely, this means that $\mathinner{|\text{nontriv}\rangle}=\mathinner{|0\rangle}$, $\mathinner{|\text{triv}\rangle}=\mathinner{|2\rangle}$,
$\mathinner{|\bot\rangle}=\sqrt{\frac{3}{4}}\mathinner{|0\rangle}+\sqrt{\frac{1}{4}}\mathinner{|2\rangle},$
[…]
$R=\mathinner{|\tilde{0}\rangle}\mathinner{\langle\tilde{0}|}-\mathinner{|\tilde{1}\rangle}\mathinner{\langle\tilde{1}|}$,
where
$\mathinner{|\tilde{0}\rangle}$ $=\frac{1}{\sqrt{2}}\mathinner{|1\rangle}-\frac{1}{2\sqrt{2}}\mathinner{|0\rangle}+\frac{\sqrt{3}}{2\sqrt{2}}\mathinner{|2\rangle}$
$\mathinner{|\tilde{1}\rangle}$ $=\frac{1}{\sqrt{2}}\mathinner{|1\rangle}+\frac{1}{2\sqrt{2}}\mathinner{|0\rangle}-\frac{\sqrt{3}}{2\sqrt{2}}\mathinner{|2\rangle}$
and thus we have determined the strategy.
[…]
3.2. Method for obtaining an upper bounds and algebraic relations
In order to prove that the above strategy is optimal, we obtain a sum-of-squares (SOS) decomposition using a method similar to that of the NPA hierarchy (NPA (08)). Let us start by defining the following sets of monomials
$A^{s}B^{t}=\{p^{x_{1}}_{a_{1}}...p^{x_{s}}_{a_{s}}q^{y_{1}}_{b_{1}}...q^{y_{t}}_{b_{t}}\mid x_{1},...,x_{s}\in X,a_{1},...,a_{s}\in A,y_{1},...,y_{t}\in Y,b_{1},...,b_{t}\in B\}$
where $s,t\in\mathbb{N}_{0}$. We denote a level of the hierarchy by
$k+A^{s_{1}}B^{t_{1}}+...+A^{s_{m}}B^{t_{m}}:=\bigcup_{s+t\leq k}A^{s}B^{t}\cup\bigcup_{i=1}^{m}A^{s_{i}}B^{t_{i}}$
where $k\in\mathbb{N}$ and $m\in\mathbb{N}_{0}$ such that $s_{i}+t_{i}>k$ for all $i\in\{1,...,m\}$.
The method can now be described as follows.
  • (1)
    Specify a level as described above, denoted $L$, and a vector of polynomials,
    $F(p^{x}_{a},q^{y}_{b})=\left(f_{1}(p^{x}_{a},q^{y}_{b}),...f_{n}(p^{x}_{a},q^{y}_{b})\right)^{T}$
    where $n\in\mathbb{N}$ and for each $i\in\{1,...,n\}$, $f_{i}(p^{x}_{a},q^{y}_{b})$ contains monomials from the set $L$.
  • (2)
    Specify an $n$-by-$n$ positive definite real matrix, $Y$, such that
    $F^{*}YF=\mu_{L}-\beta.$ (4)
    where $\mu_{L}\geq 0$.
Notice that $n$ may be strictly smaller than the cardinality of $L$. Notice also that we require $Y$ to be positive definite and not just positive semi-definite. Now, let
$S=(\mathinner{|\psi\rangle}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B},\pi_{A},\pi_{B})$
be a quantum strategy. Since $Y$ is positive definite it has a spectral decomposition
$Y=\sum_{i=1}^{n}\lambda_{i}|\varphi_{i}\rangle\langle\varphi_{i}|$
where $\lambda_{i}>0$ and the $\mathinner{|\varphi_{i}\rangle}$’s constitute an orthonormal basis of $\mathbb{C}^{n}$. Let $\tilde{F}_{i}:=\sqrt{\lambda_{i}}\mathinner{\langle\varphi_{i}|}F$ which are polynomials. Then (4) can be written as
$\sum_{i=1}^{n}\tilde{F}_{i}^{*}\tilde{F}_{i}=\mu_{L}-\beta.$
so we get
$\sum_{i=1}^{n}\pi_{A}\otimes\pi_{B}\left(\tilde{F}_{i}^{*}\tilde{F}_{i}\right)=\mu_{L}I-\pi_{A}\otimes\pi_{B}(\beta)\succeq 0.$
Thus, we obtain the upper bound
$\omega(G_{F},S)=\mathinner{\langle\psi|}\pi_{A}\otimes\pi_{B}(\beta)\mathinner{|\psi\rangle}\leq\mu_{L}.$
We can also deduce that certain algebraic relations must hold if $S$ in fact reaches the upper bound $\mu_{L}$. Suppose $\omega(G_{F},S)=\mu_{L}$. Then
$\mathinner{\langle\psi|}\pi_{A}\otimes\pi_{B}\left(\tilde{F}_{i}^{*}\tilde{F}_{i}\right)\mathinner{|\psi\rangle}=0\quad\forall i\in\{1,...,n\}$
which implies
$\mathinner{\langle\psi|}\pi_{A}\otimes\pi_{B}\left(F^{*}|\varphi_{i}\rangle\langle\varphi_{i}|F\right)\mathinner{|\psi\rangle}=0\quad\forall i\in\{1,...,n\}$
since $\lambda_{i}>0$. Now, summing over $i$ we obtain
$\mathinner{\langle\psi|}\pi_{A}\otimes\pi_{B}\left(F^{*}F\right)\mathinner{|\psi\rangle}=\sum_{i=1}^{n}\mathinner{\langle\psi|}\pi_{A}\otimes\pi_{B}\left(f_{i}^{*}f_{i}\right)\mathinner{|\psi\rangle}=\sum_{i=1}^{n}\lVert\pi_{A}\otimes\pi_{B}\left(f_{i}\right)\mathinner{|\psi\rangle}\rVert^{2}=0.$
Thus, we get
$\pi_{A}\otimes\pi_{B}\left(f_{i}\right)\mathinner{|\psi\rangle}=0\quad\forall i\in\{1,...,n\}.$
Our approach is as follows. First, we use standard NPA techniques to verify the upper bound. In the notation here, this means letting the coordinate functions $f_{i}$ in $F$ be the elements of $L$ for some specific level of hierarchy and solving the following semi-definite program
$\begin{split}\min&\text{ }\mu_{L}\\ \text{such that }&F^{*}\tilde{Y}F=\mu_{L}-\beta\\ &\tilde{Y}\succeq 0\end{split}$ (5)
From the output of this we are able to guess the relations that hold at optimality. Concretely, this is done by performing Gaussian elimination on the matrix consisting of the eigenvectors of $\tilde{Y}$ associated with nonzero eigenvalues. We then use these relations as in our polynomial basis $F$. Then we solve the following feasibility problem
$\begin{split}\min&\text{ }0\\ \text{such that }&F^{*}\tilde{Y}F=\mu_{L}-\beta\\ &\tilde{Y}-\varepsilon I\succeq 0\end{split}$ (6)
for a small $\varepsilon>0$ where $\mu_{L}$ is fixed as the output of (5). The output of (6) we then use as our matrix $Y$ in (2) above.
In the case of Feige’s game, in the standard NPA approach, using level $1+AB$ suffices to determine that $\frac{9}{16}$ is an upper bound on $\omega(G_{F})$. However, in this work, in order to achieve the desired algebraic relations, we take $L$ to be $1+AB+A^{2}+A^{3}$ and $F$ as specified in Appendix A. Moreover we are also able to round the numerical output to an exact solution, $Y$, with rational entries satisfying
$F^{*}YF=\frac{9}{16}-\beta.$
We have stored this exact solution in a Mathematica notebook in a GitHub repository SoS . Running the notebook verifies the claims made in this paper.
$3.$ The non-signalling value of Feige’s game
We have seen that the classical value of Feige’s game is $\frac{1}{2}$, and we can exceed this value with a quantum strategy that wins with probability $\frac{9}{16}$. In this section, we will study the non-signalling value of $G_{F}$ and prove $\omega_{ns}(G_{F})=\frac{2}{3}$. Recall the definition of non-signalling strategies from 2.4.
We start with the description of a non-signalling strategy that achieves the value $\frac{2}{3}$. Consider the non-signalling strategy $p$, where
[…]
We see that $p$ is a probability distribution and it holds
$\sum_{a}p(a,\perp|x,y)$ $=p(y,\perp|x,y)=\frac{1}{3},$
$\sum_{a}p(a,x|x,y)$ $=p(\perp,x|x,y)=\frac{1}{3},$
$\sum_{a}p(a,\overline{x}|x,y)$ $=p(\overline{y},\overline{x}|x,y)=\frac{1}{3},$
where $\overline{x}$ is the complement of the bit $x$. This shows $\sum_{a}p(a,b|x,y)=\sum_{a}p(a,b|x^{\prime},y)$ and similarly $\sum_{b}p(a,b|x,y)=\sum_{b}p(a,b|x,y^{\prime})$ holds. We conclude that $p$ is a non-signalling strategy. We obtain
$\omega(G_{F},p)$ $=\frac{1}{4}\sum_{x,y\in\{0,1\}}p(\perp,x|x,y)+p(y,\perp|x,y)=\frac{2}{3},$
since $p(\perp,x|x,y)=p(y,\perp|x,y)=\frac{1}{3}$.
For the upper bound of the winning probability of non-signalling strategies, consider any non-signalling strategy $p$ for $G_{F}$. We first note that
$p(\perp,x|x,y)+p(\perp,\overline{x}|\overline{x},y)+p(y,\perp|x,y)$ $\leq\sum_{a}p(a,x|x,y)+p(a,\overline{x}|\overline{x},y)+p(a,\perp|x,y)$
$=\sum_{a}p(a,x|x,y)+p(a,\overline{x}|x,y)+p(a,\perp|x,y)$
$=1,$ (7)
by the non-signalling property of $p$ and since it is a probability distribution. Similary, it holds
$p(\perp,x|x,y)+p(\overline{y},\perp|x,\overline{y})+p(y,\perp|x,y)\leq 1.$ (8)
We compute
$12\omega(G_{F},p)$ $=3\sum_{x,y\in\{0,1\}}p(\perp,x|x,y)+p(y,\perp|x,y)$
$=2\sum_{x,y\in\{0,1\}}p(\perp,x|x,y)+p(y,\perp|x,y)+\sum_{x,y\in\{0,1\}}p(\perp,\overline{x}|\overline{x},y)+p(\overline{y},\perp|x,\overline{y})$
$=\sum_{x,y\in\{0,1\}}p(\perp,x|x,y)+p(\perp,\overline{x}|\overline{x},y)+p(y,\perp|x,y)$
$\hskip 42.67912pt+\sum_{x,y\in\{0,1\}}p(\perp,x|x,y)+p(\overline{y},\perp|x,\overline{y})+p(y,\perp|x,y)$
$\leq 8,$
by rearranging terms and using Sections 3.3 and 8 for the inequality. This yields $\omega(G_{F},p)\leq\frac{2}{3}$. Summarizing, we obtain $\omega_{ns}(G_{F})=\frac{2}{3}$.
4. Feige’s game is a self-test
The goal of this section is to show that $\tilde{S}_{F}$ given in Section 3.1 is in fact the unique optimal quantum strategy of $G_{F}$ in the sense made precise in the following definition.
Definition 4.1.
Let $G=(X,Y,A,B,\mu,V)$ be a nonlocal game, and let
$\tilde{S}=(\mathinner{|\tilde{\psi}\rangle}\in\tilde{\mathcal{H}}_{A}\otimes\tilde{\mathcal{H}}_{B},\{\tilde{P}^{x}_{a}\},\{\tilde{Q}^{y}_{b}\})$
be an optimal quantum strategy. We say that $G$ robustly self-tests $\tilde{S}$ if there exists a function
$\delta:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0},\qquad\delta(\epsilon)\rightarrow 0\text{ as }\epsilon\rightarrow 0,$
such that for any $\epsilon\geq 0$ and any strategy
$S=(\mathinner{|\psi\rangle}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B},\{P^{x}_{a}\},\{Q^{y}_{b}\}),$
with $\omega(G,S)=\omega_{q}(G)-\epsilon$, there exist isometries
$V_{A}:\mathcal{H}_{A}\rightarrow\tilde{\mathcal{H}}_{A}\otimes\mathcal{K}_{A}\text{, }V_{B}:\mathcal{H}_{B}\rightarrow\tilde{\mathcal{H}}_{B}\otimes\mathcal{K}_{B}$
and a unit vector $\mathinner{|\kappa\rangle}\in\mathcal{K}_{A}\otimes\mathcal{K}_{B}$ such that
$\lVert V_{A}\otimes V_{B}\left(P^{x}_{a}\otimes Q^{y}_{b}\mathinner{|\psi\rangle}\right)-\left(\tilde{P}^{x}_{a}\otimes\tilde{Q}^{y}_{b}\mathinner{|\tilde{\psi}\rangle}\right)\otimes\mathinner{|\kappa\rangle}\rVert\leq\delta(\epsilon)$ (9)
for all $x,y,a,b$.
Recall the following definition from (Zha, 24, Definitions 6.1 and 7.1).
Definition 4.2.
Let $\Gamma=\{\gamma^{y}_{b}:b\in B,y\in Y\}$ and $\mathcal{R}$ be two sets of *-polynomials in $\mathscr{A}_{PVM}^{X,A}$, where every $\gamma^{y}_{b}$ is self-adjoint. We say $(\Gamma,\mathcal{R})$ is a determining pair for a nonlocal game $G=(X,Y,A,B,\mu,V)$ if the following conditions hold.
  1. (1)
    For every $y\in Y$, $\{q_{\mathcal{R}}(\gamma^{y}_{b}):b\in B\}$ is a PVM in the quotient $\mathscr{A}_{PVM}^{X,A}/\langle\mathcal{R}\rangle$, where $\langle\mathcal{R}\rangle$ is the two-sided closed ideal generated by $\mathcal{R}$ and
    $q_{\mathcal{R}}:\mathscr{A}_{PVM}^{X,A}\rightarrow\mathscr{A}_{PVM}^{X,A}/\langle\mathcal{R}\rangle$
    is the quotient map.
  2. (2)
    A strategy $\tilde{S}=(\mathinner{|\tilde{\psi}\rangle},\tilde{\pi}_{A},\tilde{\pi}_{B})$ is optimal if and only if
    1. (a)
      $I_{A}\otimes\tilde{\pi}_{B}(q^{y}_{b})\mathinner{|\tilde{\psi}\rangle}=\tilde{\pi}_{A}(\gamma^{y}_{b})\otimes I_{B}\mathinner{|\psi\rangle}$ for all $b,y$, and
    2. (b)
      the linear functional $\tau$ on $\mathscr{A}_{PVM}^{X,A}$ defined by
      $\tau(\alpha):=\mathinner{\langle\tilde{\psi}|}\tilde{\pi}_{A}(\alpha)\otimes I_{B}\mathinner{|\tilde{\psi}\rangle}$
      is a tracial state satisfying $\tau(r^{*}r)=0$ for all $r\in\mathcal{R}$.
In this case, we define the game algebra of $G$ as
$C^{*}(G):=\mathscr{A}_{PVM}^{X,A}/\langle\mathcal{R}\rangle.$
We say that a determining pair $(\Gamma,\mathcal{R})$ is $\delta$-robust for some function
$\delta:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0},\qquad\delta(\epsilon)\rightarrow 0\text{ as }\epsilon\rightarrow 0$
if for any $\epsilon\geq 0$ and any quantum strategy $S=(\mathinner{|\psi\rangle},\pi_{A},\pi_{B})$ with $\omega(G,S)=w_{q}(G)-\epsilon$,
  1. (3)
    $\lVert I_{A}\otimes\pi_{B}(q^{y}_{b})\mathinner{|\psi\rangle}-\pi_{A}(\gamma^{y}_{b})\otimes I_{B}\mathinner{|\psi\rangle}\rVert\leq\delta(\epsilon)$ for all $b,y$, and
  2. (4)
    $\lVert\pi_{A}(r)\otimes I_{B}\mathinner{|\psi\rangle}\rVert\leq\delta(\epsilon)$ for all $r\in\mathcal{R}$.
The optimal strategy $\tilde{S}$ in 4.2 is sometimes said to be Alice-determined: Bob’s PVMs $\{q^{y}_{b}\}$ are determined by some PMVs $\Gamma=\{\gamma^{y}_{b}\}$ on Alice’s side (see (2a)), and Alice’s PVMs satisfy the algebraic relations $\mathcal{R}$ (see (2b)). A nearly optimal strategy is then Alice-robustly-determined, meaning that the equalities in (2a) and (2b) still hold approximately. In particular, Alice’s PVMs in an optimal (resp. nearly optimal) strategy define a representation (resp. an approximate representation) of the associated game algebra $C^{*}(G)$.
Thus, to prove $G$ has a unique optimal strategy in the sense of definition 4.1, it suffices to show that $C^{*}(G)$ has a unique irreducible representation, and the robustness of this self-test then follows from the robustness of the determining pair.
Theorem 4.3 (Corollary 7.11 of Zha (24)).
Let $G$ be a nonlocal game with a $\delta$-robust determining pair $(\Gamma,\mathcal{R})$. If the associated game algebra $C^{*}(G)$ is a full matrix algebra, then $G$ is a robust self-test.
Within this framework, to prove that $G_{F}$ robustly self-tests $\tilde{S}_{F}$, we need to show that
  1. (A)
    $G_{F}$ has a robust determining pair $(\Gamma,\mathcal{R})$, and
  2. (B)
    the associated game algebra $C^{*}(G_{F})$ has a unique irreducible representation.
Our approach for Task (A), constructing a determining pair and showing its robustness, is based on a sum-of-squares (SOS) decomposition of the game polynomial of $G_{F}$ as detailed in Section 3.2. We get the following algebraic structure of optimal strategies for $G_{F}$.
Lemma $4.$
Let $\nu:=\begin{bmatrix}\frac{9}{16}&\frac{1}{16}&\frac{3}{8}\\ \frac{1}{16}&\frac{9}{16}&\frac{3}{8}\\ \frac{3}{8}&\frac{3}{8}&\frac{1}{4}\end{bmatrix}$, with the first, second, and third column/row labelled by $0,1$, and $\bot$, respectively. Let
$\Gamma:=\{\gamma^{y}_{b}:b\in\{0,1,\bot\},y\in\{0,1\}\}\subseteq\mathscr{A}_{PVM}^{\{0,1\},\{0,1,\bot\}}$
where
$\gamma^{0}_{0}:$ $=-\tfrac{3}{4}\{p^{0}_{1},p^{1}_{0}\}-\tfrac{5}{4}\{p^{0}_{1},p^{1}_{1}\}-\tfrac{5}{16}p^{0}_{0}+\tfrac{3}{8}p^{1}_{0}+\tfrac{9}{16}p^{0}_{1}+\tfrac{15}{8}p^{1}_{1},$
$\gamma^{0}_{1}:$ $=-\tfrac{3}{4}\{p^{0}_{1},p^{1}_{0}\}-\tfrac{5}{4}\{p^{0}_{1},p^{1}_{1}\}+\tfrac{3}{16}p^{0}_{0}-\tfrac{1}{8}p^{1}_{0}+\tfrac{33}{16}p^{0}_{1}+\tfrac{3}{8}p^{1}_{1},$
$\gamma^{0}_{\bot}:$ $=1-\gamma^{0}_{0}-\gamma^{0}_{1},$
$\gamma^{1}_{0}:$ $=-\tfrac{3}{4}\{p^{0}_{1},p^{1}_{0}\}-\tfrac{5}{4}\{p^{0}_{1},p^{1}_{1}\}-\tfrac{9}{16}p^{0}_{0}+\tfrac{9}{8}p^{1}_{0}+\tfrac{13}{16}p^{0}_{1}+\tfrac{9}{8}p^{1}_{1},$
$\gamma^{1}_{1}:$ $=-\tfrac{3}{4}\{p^{0}_{1},p^{1}_{0}\}-\tfrac{5}{4}\{p^{0}_{1},p^{1}_{1}\}+\tfrac{15}{16}p^{0}_{0}-\tfrac{3}{8}p^{1}_{0}+\tfrac{21}{16}p^{0}_{1}+\tfrac{5}{8}p^{1}_{1},$
$\gamma^{1}_{\bot}:$ $=1-\gamma^{1}_{0}-\gamma^{1}_{1}.$
Here $\{p,q\}:=pq+qp$ is the anti-commutator. Let
$\mathcal{R}:=\{p^{x}_{a}p^{1-x}_{a^{\prime}}p^{x}_{a}-\nu_{aa^{\prime}}p^{x}_{a}:a,a^{\prime}\in\{0,1,\bot\},x\in\{0,1\}\}\subseteq\mathscr{A}_{PVM}^{\{0,1\},\{0,1,\bot\}}.$
Then $(\Gamma,\mathcal{R})$ is an $O(\sqrt{\epsilon})$-robust determining pair of $G_{F}$.
Proof.
It is not hard to verify that $\gamma^{y}_{b}$’s are self-adjoint such that $\sum_{b}\gamma^{y}_{b}=1$ for $y\in\{0,1\}$, and $(\gamma^{y}_{b})^{2}-\gamma^{y}_{b}\in\langle\mathcal{R}\rangle$ for all $y,b$. So for every $y$, $\{\gamma^{y}_{b}:b\in\{0,1,\bot\}\}$ is a PVM in $\mathscr{A}_{PVM}^{\{0,1\},\{0,1,\bot\}}/\langle\mathcal{R}\rangle.$ This proves Hypothesis (1) in 4.2.
Let $\beta\in\mathscr{A}_{PVM}^{X,A}\otimes\mathscr{A}_{PVM}^{Y,B}$ be the game polynomial of $G_{F}$, and let
$F=F(p_{a}^{x},q_{b}^{y})=\left(f_{1}(p_{a}^{x},q_{b}^{y}),...,f_{32}(p_{a}^{x},q_{b}^{y})\right)^{T}$
be the vector of 32 polynomials in $p_{a}^{x}$ and $q_{b}^{y}$ given in Appendix A. Let $Y$ be the $32\times 32$ positive definite matrix given in SoS , then as stated in Section 3.2, using a computer, one can check that
$F^{*}YF=\tfrac{9}{16}-\beta.$
As in Section 3.2 we write $Y\in M_{32}(\mathbb{C})$ in its spectral decomposition
$Y=\sum_{i=1}^{32}\lambda_{i}|\varphi_{i}\rangle\langle\varphi_{i}|,$
where $\lambda_{i}>0$ for all $i$ and $\{\mathinner{|\varphi_{i}\rangle}:1\leq i\leq 32\}$ is an orthonormal basis of $\mathbb{C}^{32}$. For any quantum strategy $S=(\mathinner{|\psi\rangle},\pi_{A},\pi_{B})$, let
$\hat{F}=\big((\pi_{A}\otimes\pi_{B})(f_{1}),\ldots,(\pi_{A}\otimes\pi_{B})(f_{32})\big)^{T},$
and let $\epsilon=\frac{9}{16}-\omega(G_{F},S)$. It follows that
$\epsilon$ $=\mathinner{\langle\psi|}(\pi_{A}\otimes\pi_{B})\left(\frac{9}{16}-\beta\right)\mathinner{|\psi\rangle}=\mathinner{\langle\psi|}\hat{F}^{*}Y\hat{F}\mathinner{|\psi\rangle}$
$=\sum_{i=1}^{32}\lambda_{i}\mathinner{\langle\psi|}\hat{F}^{*}\mathinner{|\varphi_{i}\rangle}\mathinner{\langle\varphi_{i}|}\hat{F}\mathinner{|\psi\rangle}=\sum_{i=1}^{32}\lambda_{i}\lvert\mathinner{\langle\varphi_{i}|}\hat{F}\mathinner{|\psi\rangle}\rvert^{2}.$
Since each $\lambda_{i}>0$, we obtain that
$\lvert\mathinner{\langle\varphi_{i}|}\hat{F}\mathinner{|\psi\rangle}\rvert^{2}\leq O(\epsilon)$
for all $1\leq i\leq 32$. Hence
$\sum_{i=1}^{32}\lVert(\pi_{A}\otimes\pi_{B}(f_{i})\mathinner{|\psi\rangle}\rVert^{2}=\mathinner{\langle\psi|}\hat{F}^{*}\hat{F}\mathinner{|\psi\rangle}=\sum_{i=1}^{32}\lvert\mathinner{\langle\varphi_{i}|}\hat{F}\mathinner{|\psi\rangle}\rvert^{2}\leq O(\epsilon).$
In particular,
$\lVert(\pi_{A}\otimes\pi_{B})(f_{i})\mathinner{|\psi\rangle}\rVert\leq O(\sqrt{\epsilon})$
for every $1\leq i\leq 32$. We further observe that
$1\otimes q^{y}_{b}-\gamma^{y}_{b}\otimes 1,r\otimes 1\in\operatorname{span}\{f_{i}:1\leq i\leq 32\}$
for all $b,y$ and all $r\in\mathcal{R}$. We conclude that
$\lVert I_{A}\otimes\pi_{B}(q^{y}_{b})\mathinner{|\psi\rangle}-\pi_{A}(\gamma^{y}_{b})\otimes I_{B}\mathinner{|\psi\rangle}\rVert\leq O(\sqrt{\epsilon})$ (10)
for all $b,y$, and
$\lVert\pi_{A}(r)\otimes I_{B}\mathinner{|\psi\rangle}\rVert\leq O(\sqrt{\epsilon})$ (11)
for all $r\in\mathcal{R}$. This proves Hypotheses (3) and (4) in 4.2 with $\delta(\epsilon)=O(\sqrt{\epsilon})$.
Now suppose $S$ is optimal for $G_{F}$, then $\epsilon=0$, so the “only if” part of Hypothesis (2a) follows from Equation 10. By Equation 11, the linear functional $\tau$ on $\mathscr{A}_{PVM}^{\{0,1\},\{0,1,\bot\}}$ defined by
$\tau(\alpha):=\mathinner{\langle\psi|}\pi_{A}(\alpha)\otimes I_{B}\mathinner{|\psi\rangle}$
is a state satisfying $\tau(r^{*}r)=0$ for all $r\in\mathcal{R}$. By symmetry, for every $p^{x}_{a}$ there exists self-adjoint $\eta^{x}_{a}\in\mathscr{A}_{PVM}^{Y,B}$ such that $\pi_{A}(p^{x}_{a})\otimes I_{B}\mathinner{|\psi\rangle}=I_{A}\otimes\pi_{B}(\eta^{x}_{a})\mathinner{|\psi\rangle}$. So for any monomials $w_{1}=p^{x_{1}}_{a_{1}}\cdots p^{x_{k}}_{a_{k}}$ and $w_{2}=p^{x_{k+1}}_{a_{k+1}}\cdots p^{x_{n}}_{a_{n}}$, we have
$\mathinner{\langle\psi|}\pi_{A}(w_{2}w_{1})\otimes I_{B}\mathinner{|\psi\rangle}$ $=\mathinner{\langle\psi|}\pi_{A}(w_{2})\pi_{A}(p^{x_{1}}_{a_{1}}\cdots p^{x_{k-1}}_{a_{k-1}})\otimes\pi_{B}(\eta^{x_{k}}_{a_{k}})\mathinner{|\psi\rangle}$
$=\mathinner{\langle\psi|}\pi_{A}(p^{x_{k}}_{a_{k}})\pi_{A}(w_{2})\pi_{A}(p^{x_{1}}_{a_{1}}\cdots p^{x_{k-1}}_{a_{k-1}})\otimes I_{B}\mathinner{|\psi\rangle}$
$=\cdots$
$=\mathinner{\langle\psi|}\pi_{A}(p^{x_{1}}_{a_{1}}\cdots p^{x_{k}}_{a_{k}})\pi_{A}(w_{2})\otimes I_{B}\mathinner{|\psi\rangle}$
$=\mathinner{\langle\psi|}\pi_{A}(w_{1}w_{2})\otimes I_{B}\mathinner{|\psi\rangle}.$
Hence $\tau$ is tracial. This proves the “only if” part of (2b).
To see the “if” part of Hypothesis (2), suppose $\tilde{S}=(\mathinner{|\tilde{\psi}\rangle},\tilde{\pi}_{A},\tilde{\pi}_{B})$ is a strategy satisfying (2a) and (2b). Then $\mathinner{\langle\tilde{\psi}|}(\tilde{\pi}_{A}\otimes\tilde{\pi}_{B})(\beta)\mathinner{|\tilde{\psi}\rangle}=\frac{9}{16}$. Hence $\tilde{S}$ is optimal. This completes the proof. ∎
For Task (B), proving the game algebra $C^{*}(G_{F})$ has a unique irreducible representation, we first note that by 4.4, $C^{*}(G)=\mathscr{A}_{PVM}^{\{0,1\},\{0,1,\bot\}}/\langle\mathcal{R}\rangle$ is isometric to $\mathcal{A}_{\nu}$, the universal $C^{*}$-algebra generated by $E_{i},F_{i},1\leq i\leq 3$, subject to the relations
  1. (i)
    $E_{i}=E_{i}^{*}$, $F_{i}^{*}=F_{i}$ for all $i$,
  2. (ii)
    $\sum_{i=1}^{3}E_{i}=\sum_{i=1}^{3}F_{i}=1$, and
  3. (iii)
    $E_{i}F_{j}E_{i}=\nu_{ij}E_{i}$, $F_{i}E_{j}F_{i}=\nu_{ij}F_{i}$ for all $i,j$,
where $\nu$ is the symmetric stochastic matrix defined in $4.$ In fact, for any $n\times n$ symmetric stochastic matrix $\nu$, we can define $\mathcal{A}_{\nu}$ analogously, where generators $\{E_{i}\}_{i=1}^{n}$ and $\{F_{i}\}_{i=1}^{n}$ are PVMs satisfying $E_{i}F_{j}E_{i}=\nu_{ij}E_{i}$ and $F_{i}E_{j}F_{i}=\nu_{ij}F_{i}$. In Appendix B, we establish the representation theory for those $\mathcal{A}_{\nu}$’s. Interestingly, we show that representations for $\mathcal{A}_{\nu}$ can be seen as a generalization of mutually unbiased measurements (MUMs, see e.g. TFR+ (21)). We generalize the results in TFR+ (21) and give a characterization of irreducible representations of $\mathcal{A}_{\nu}$. In particular, in B.9, we show that for $2\times 2$ and $3\times 3$ symmetric stochastic matrices $\nu$, every $\mathcal{A}_{\nu}$ has a unique irreducible representation up to unitary equivalence and relabeling generators. For the matrix $\nu$ defined in 4.4, we further show that any allowed relabeling still results in a unitarily equivalent representation.
Lemma 4.5.
The game algebra $C^{*}(G_{F})$ associated with $G_{F}$ has a unique irreducible representation up to unitary equivalence. Moreover, $G_{F}$ is isomorphic to the full matrix algebra $M_{3}$.
We prove this lemma in Appendix B.
Theorem 4.6.
The game $G_{F}$ robustly self-tests the strategy $\tilde{S}_{F}$.
Proof.
Let $(\Gamma,\mathcal{R})$ be the $O(\sqrt{\epsilon})$-robust determining pair of $G_{F}$ defined in $4.$ By 4.5, the associated game algebra $C^{*}(G)=\mathscr{A}_{PVM}^{\{0,1\},\{0,1,\bot\}}/\langle\mathcal{R}\rangle$ is a full matrix algebra. It follows from 4.3 that $G_{F}$ is a robust self-test.
5. Feige’s game as $(G_{1}\lor G_{2})$-game
Let $G_{1}$ and $G_{2}$ be nonlocal games. We describe the $(G_{1}\lor G_{2})$-game as introduced in MS (23). The referee sends Alice and Bob a pair of questions $(x_{1},x_{2})$ and $(y_{1},y_{2})$, respectively, where $x_{i},y_{i}$ are questions in $G_{i}$, $i\in\{1,2\}$. Each of them chooses one of the questions they received and responds with an answer from the corresponding game. To win the game, two conditions have to be fulfilled:
  • (1)
    Alice and Bob have to give answers from the same game,
  • (2)
    their answers have to win the corresponding game.
More formally, suppose the nonlocal games $G_{1}$ and $G_{2}$ have input sets $X_{i}$, $Y_{i}$, output sets $A_{i}$, $B_{i}$, verification functions $V_{G_{i}}$ and probability distributions $\pi_{i}$ on $X_{i}\times Y_{i}$ for $i=1,2$. Then the $(G_{1}\lor G_{2})$-game has input sets $X_{1}\times X_{2}$, $Y_{1}\times Y_{2}$, output sets $A_{1}\dot{\cup}A_{2}$, $B_{1}\dot{\cup}B_{2}$ and verification function
[…]
For the probability distribution, we take $\pi=\pi_{1}\times\pi_{2}$ on $(X_{1}\times X_{2})\times(Y_{1}\times Y_{2})$, i.e. $\pi((x_{1},x_{2}),(y_{1},y_{2}))=\pi_{1}(x_{1},y_{1})\pi_{2}(x_{2},y_{2})$. In this article, we assume that the probability distributions of $G_{1}$ and $G_{2}$ are uniform, so this will also be the case for the $(G_{1}\lor G_{2})$-game.
We will now show that there exist $G_{1}$ and $G_{2}$ such that $G_{F}=G_{1}\lor G_{2}$, where $G_{F}$ denotes the Feige game. We will see $\frac{9}{16}=\omega_{q}(G_{F})>\mathrm{max}\{\omega_{q}(G_{1}),\omega_{q}(G_{2})\}=\frac{1}{2}$. This is in contrast the case of perfect strategies (MS, 23, Lemma 4.3), where $\omega_{q}(G_{1}\lor G_{2})=1$ implies $\omega_{q}(G_{i})=1$ if the players give answers for $G_{i}$ a perfect strategy of the ”or”-game.
Consider the games $G_{1}$ and $G_{2}$, where
$X_{1}=\{*\},\quad Y_{1}=\{0,1\},\quad A_{1}=\{0,1\},\quad B_{1}=\{\perp\},\quad\pi(x,y)=\frac{1}{2}$ (12)
$X_{2}=\{0,1\},\quad Y_{2}=\{*\},\quad A_{2}=\{\perp\},\quad B_{2}=\{0,1\},\quad\pi(x,y)=\frac{1}{2},$ (13)
$V_{1}(a,\perp,*,y)=\begin{cases}1\text{ if }a=y\\ 0\text{ otherwise,}\end{cases}V_{2}(\perp,b,x,*)=\begin{cases}1\text{ if }b=x\\ 0\text{ otherwise.}\end{cases}$ (14)
In the game $G_{1}$, Alice guesses the bit that Bob received and their roles are reversed in the game $G_{2}$.
Lemma 5.1.
It holds $\omega_{q}(G_{i})=\frac{1}{2}$ for $i=1,2$.
Proof.
We prove it for $G_{1}$, the proof for $G_{2}$ is the same. Note that it holds $B^{y}_{\perp}=\mathrm{Id}$ for each Bob POVM in any quantum strategy $S$, since he is only allowed to answer with $\perp$. Thus, the winning probability of any quantum strategy for $G_{1}$ is given by
$\omega_{q}(S,G_{1})=\frac{1}{2}\mathinner{\langle\psi|}A^{*}_{0}\otimes B^{0}_{\perp}+A^{*}_{1}\otimes B^{1}_{\perp}\mathinner{|\psi\rangle}=\frac{1}{2}\mathinner{\langle\psi|}(A^{*}_{0}+A^{*}_{1})\otimes\mathrm{Id}\mathinner{|\psi\rangle}=\frac{1}{2},$
where we used $A^{*}_{0}+A^{*}_{1}=\mathrm{Id}$ in the last step. This completes the proof. ∎
Theorem 5.2.
It holds $G_{F}=G_{1}\lor G_{2}$. We therefore have
$\omega_{q}(G_{F})=\frac{9}{16}>\frac{1}{2}=\mathrm{max}\{\omega_{q}(G_{1}),\omega_{q}(G_{2})\}.$
Proof.
From Equations 12 and 13, we deduce
$X_{G_{1}\lor G_{2}}$ $=\{(*,0),(*,1)\},$ $Y_{G_{1}\lor G_{2}}=\{(0,*),(1,*)\},$
$A_{G_{1}\lor G_{2}}$ $=\{0,1,\perp\},$ $B_{G_{1}\lor G_{2}}=\{0,1,\perp\}.$
Now, relabeling the sets $\{(*,0),(*,1)\}$ and $\{(0,*),(1,*)\}$ by $\{0,1\}$, we obtain that $G_{1}\lor G_{2}$ has the same question and answer sets as the Feige game. It is straightforward to see that the verification functions coincide. ∎
6. Feige’s game and parallel repetition
In this section, we look at parallel repetition of Feige’s game. From Table 1, we see that Feige’s game has interesting behaviour with respect to parallel repetition. In the table, the classical, quantum and non-signalling values agree in the case of an even number of repetitions, but this is not the case for odd numbers. As we saw previously, all values are different for the original game. Furthermore, we get that at least the classical and non-signalling value are different for $n=3$, despite this not being the case for $n=2$. We do not know if there is quantum advantage for $n=3$. Note that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition performs worse than the optimal classical strategy for $n=3$, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$.
We first recall the definition of parallel repetition. In the following, we define the product of two games, which then leads to the definition of the parallel repetition of a game.
Definition 6.1.
Suppose
$\begin{split}G_{1}=(X_{1},Y_{1},A_{1},B_{1},\pi_{1},V_{1}),\quad G_{2}=(X_{2},Y_{2},A_{2},B_{2},\pi_{2},V_{2})\end{split}$
are nonlocal games. We define their product as the game
$G_{1}\times G_{2}=\big(X_{1}\times X_{2},Y_{1}\times Y_{2},A_{1}\times A_{2},B_{1}\times B_{2},\pi_{1}\times\pi_{2},V_{1}\times V_{2}\big)$
Where Alice and Bob receive pairs of questions $(x_{1},x_{2})\in X_{1}\times X_{2}$ and $(y_{1},y_{2})\in Y_{1}\times Y_{2}$ respectively, and provide answers $(a_{1},a_{2})\in A_{1}\times A_{2}$ and $(b_{1},b_{2})\in B_{1}\times B_{2}$ respectively. The prior distribution is given by
$\pi_{1}\times\pi_{2}\left((x_{1},x_{2}),(y_{1},y_{2})\right):=\pi_{1}(x_{1},y_{1})\pi_{2}(x_{2},y_{2}).$
and $V_{1}\times V_{2}$ likewise
$V_{1}\times V_{2}\Big((a_{1},a_{2}),(b_{1},b_{2})|(x_{1},x_{2}),(y_{1},y_{2})\Big)=V_{1}(a_{1},b_{1}|x_{1},y_{1})V_{2}(a_{2},b_{2}|x_{2},y_{2})$
In words, Alice and Bob play the two games independently in parallel and they win if and only if they win both of them simultaneously.
Definition 6.2.
Let $G=(X,Y,A,B,\pi,V)$ be a nonlocal game. The game $G^{\times n}$ is called the $n$-fold parallel repetition of $G$.
6.1. Classical strategies for the $2m$-fold parallel repetition
The following strategy was already described by Feige Fei (91) for the game $G_{F}^{\times 2}$. Receiving question $(x_{1},x_{2})$ and $(y_{1},y_{2})$, Alice and Bob answer
$a_{1}$ $=\perp,$ $b_{1}=y_{2},$
$a_{2}$ $=x_{1},$ $b_{2}=\perp,$
respectively. We see that the players win if and only if $x_{1}=y_{2}$, which happens with probability $\frac{1}{2}$. This may be surprising, since the classical winning probability of the 22-fold parallel repetition does not decrease when compared with the original game $G_{F}$.
For the $2m$-fold parallel repetition $G_{F}^{\times 2m}$, we obtain a classical strategy with winning probability $\frac{1}{2^{m}}=\frac{1}{2^{n/2}}$ by repeating the strategy above $m$ times. More precisely, given $(x_{1},\dots,x_{2m})$ and $(y_{1},\dots,y_{2m})$, the players answer $(a_{1},\dots,a_{2m})$ and $(b_{1},\dots,b_{2m})$ with
$a_{2i-1}$ $=\perp,$ $b_{2i-1}=y_{2i},$
$a_{2i}$ $=x_{2i-1},$ $b_{2i}=\perp,$
for $1\leq i\leq m$. We will see in the next section, that this probability coincides with the upper bound on the non-signalling value of $G_{F}^{\times 2m}$.
6.2. The non-signalling upper bound for $2m$-fold parallel repetition
Let $n=2m$. Consider a set $I\subseteq[n]:=\{1,\dots,n\}$ and define $a_{I}(\mathbf{y})\in\{0,1,\perp\}^{n},b_{I}(\mathbf{x})\in\{0,1,\perp\}^{n}$ for each $\mathbf{x}$, $\mathbf{y}\in\{0,1\}^{n}$ by
$(a_{I}(\mathbf{y}))_{i}=\begin{cases}\perp,i\in I\\ y_{i},i\notin I\end{cases}\quad\text{and}\quad(b_{I}(\mathbf{x}))_{i}=\begin{cases}\perp,i\notin I\\ x_{i},i\in I.\end{cases}$
Then, we can write
$\omega(G_{F}^{\times n},p)=\frac{1}{2^{4m}}\sum_{I\subseteq[n]}\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$
for the winning probability of the $2m$-fold parallel repetition. The goal of this subsection is to show $\omega(G_{F}^{\times n},p)\leq\frac{1}{2^{m}}=\frac{1}{2^{n/2}}$, which is equivalent to proving
$\sum_{I\subseteq[n]}\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq 2^{3m}.$
For this, we first show the following lemma. We define $\mathbf{x}|_{I}\in\{0,1\}^{|I|}$ to be the restriction of $\mathbf{x}$ to a subset $I\subseteq[n]$. Furthermore, we write $|I|$ for the cardinality of a set II and denote the complement by $I^{c}=[n]\setminus I$.
Lemma 6.3.
Let $J\subseteq\{0,1\}^{|I^{c}|}$. It holds
$\sum_{\mathbf{x}|_{I^{c}}\in J}\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq|J|\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{z},\mathbf{y})$
for any $\mathbf{z}\in\{0,1\}^{n}$. Similarly, for $K\subseteq\{0,1\}^{|I|}$, it holds
$\sum_{\mathbf{y}|_{I}\in K}\sum_{\mathbf{y}|_{I}\in\{0,1\}^{|I^{c}|}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq|K|\sum_{\mathbf{y}|_{I^{c}}\in\{0,1\}^{|I^{c}|}}\sum_{\mathbf{b}}p(a_{I}(\mathbf{y}),\mathbf{b}|\mathbf{x},\mathbf{z})$
for any $\mathbf{z}\in\{0,1\}^{n}$.
Proof.
We prove the first inequality, the second follows similarly. For fixed $\mathbf{x}|_{I^{c}}$, we have
$\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{z},\mathbf{y})$ (15)
because of $p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$ and the non-signalling property of pp. Since $b_{I}(\mathbf{v})=b_{I}(\mathbf{w})$ if and only if $\mathbf{v}|_{I}=\mathbf{w}|_{I}$, we see that the right hand side of Equation 15 is the same for each $\mathbf{x}|_{I^{c}}\in J$. This yields the assertion. ∎
The next lemma is concerned with sets that have exactly $m$ elements.
Lemma 6.4.
Consider a subset $I\subseteq[n]$ with $m$ elements. Then
$\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq 2^{m-1}\Bigg(\sum_{\mathbf{y},\mathbf{a}}$ $\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{z},\mathbf{y})$
$+\sum_{\mathbf{x},\mathbf{b}}\sum_{\mathbf{y}|_{I^{c}}\in\{0,1\}^{|I^{c}|}}p(a_{I}(\mathbf{y}),\mathbf{b}|\mathbf{x},\mathbf{z})\Bigg)$
Proof.
Since $n=2m$, both sets II and $I^{c}$ have $m$ elements. Therefore, we can choose pairs $(i_{k},j_{k})$ for $1\leq k\leq m$ with $i_{k}\in I,j_{k}\in I^{c}$ such that $\{i_{k},j_{k}|1\leq k\leq m\}=[n]$. It holds
$\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$ $=\sum_{l\in\{0,1\}^{m}}\sum_{\mathbf{x},\mathbf{y}}\sum_{x_{i_{k}}\oplus y_{j_{k}}=l_{k}\text{ for all }k}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$
$=\sum_{\begin{subarray}{c}l\in\{0,1\}^{m}\\ \sum_{k}l_{k}=0\text{ mod }2\end{subarray}}\sum_{\mathbf{x},\mathbf{y}}\sum_{x_{i_{k}}\oplus y_{j_{k}}=l_{k}\text{ for all }k}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$
$\quad+\sum_{\begin{subarray}{c}l\in\{0,1\}^{m}\\ \sum_{k}l_{k}=1\text{ mod }2\end{subarray}}\sum_{\mathbf{x},\mathbf{y}}\sum_{x_{i_{k}}\oplus y_{j_{k}}=l_{k}\text{ for all }k}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y}).$ (16)
Now, 6.3 yields
$\sum_{\begin{subarray}{c}l\in\{0,1\}^{m}\\ \sum_{k}l_{k}=0\text{ mod }2\end{subarray}}\sum_{\mathbf{x},\mathbf{y}}\sum_{\begin{subarray}{c}x_{i_{k}}\oplus y_{j_{k}}=l_{k}\\ \text{ for all }k\end{subarray}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq 2^{m-1}\sum_{\mathbf{y},\mathbf{a}}\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{z},\mathbf{y})$
since the set $\{l\in\{0,1\}^{m}|\sum_{k}l_{k}=0\text{ mod }2\}$ has $2^{m-1}$ elements. In the same way, we obtain
$\sum_{\begin{subarray}{c}l\in\{0,1\}^{m}\\ \sum_{k}l_{k}=1\text{ mod }2\end{subarray}}\sum_{\mathbf{x},\mathbf{y}}\sum_{\begin{subarray}{c}x_{i_{k}}\oplus y_{j_{k}}=l_{k}\\ \text{ for all }k\end{subarray}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq 2^{m-1}\sum_{\mathbf{x},\mathbf{b}}\sum_{\mathbf{y}|_{I^{c}}\in\{0,1\}^{|I^{c}|}}p(a_{I}(\mathbf{y}),\mathbf{b}|\mathbf{x},\mathbf{z}).$
Together with Equation 16 this yields the result. ∎
Now, we can prove the main theorem of this subsection.
Theorem 6.5.
It holds $\omega_{ns}(G_{F}^{\times n})\leq\frac{1}{2^{n/2}}$ for $n$ even.
Proof.
We will prove
$2^{4m}\omega(G_{F}^{\times n},p)=\sum_{I\subseteq[n]}\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq 2^{3m}$
for any nonsignalling strategy pp. From 6.3, we obtain for $|I|>m$ that
$\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$ $=\sum_{\mathbf{y}}\sum_{\mathbf{x}|_{I^{c}}\in\{0,1\}^{|I|^{c}}}\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$
$\leq 2^{m-1}\sum_{\mathbf{y}}\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{z},\mathbf{y}),$ (17)
since $|\{0,1\}^{|I|^{c}}|\leq 2^{m-1}$. Similarly, we get
$\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq 2^{m-1}\sum_{\mathbf{x}}\sum_{\mathbf{y}|_{I^{c}}\in\{0,1\}^{|I^{c}|}}\sum_{\mathbf{b}}p(a_{I}(\mathbf{y}),\mathbf{b}|\mathbf{x},\mathbf{z})$ (18)
for $|I|<m$. Together with 6.4, Equations 17 and 18 yield
$\sum_{I\subseteq[n]}\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq 2^{m-1}\Bigg($ $\sum_{I,|I|\geq m}\sum_{\mathbf{y}}\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{z},\mathbf{y})$
$+\sum_{I,|I|\leq m}\sum_{\mathbf{x}}\sum_{\mathbf{y}|_{I^{c}}\in\{0,1\}^{|I^{c}|}}\sum_{\mathbf{b}}p(a_{I}(\mathbf{y}),\mathbf{b}|\mathbf{x},\mathbf{z})\Bigg)$ (19)
Note that we have
$\sum_{I,|I|\geq m}\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{z},\mathbf{y})\leq\sum_{\mathbf{a},\mathbf{b}}p(\mathbf{a},\mathbf{b})|\mathbf{z},\mathbf{y})=1$
since $b_{I_{1}}(\mathbf{v})\neq b_{I_{2}}(\mathbf{w})$ for $I_{1}\neq I_{2}$ and any $\mathbf{v},\mathbf{w}$. Similarly, we obtain
$\sum_{I,|I|\leq m}\sum_{\mathbf{x}}\sum_{\mathbf{y}|_{I^{c}}\in\{0,1\}^{|I^{c}|}}\sum_{\mathbf{b}}p(a_{I}(\mathbf{y}),\mathbf{b}|\mathbf{x},\mathbf{z})\leq 1.$
From Equation 19, we then deduce
$\sum_{I\subseteq[n]}\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq 2^{m-1}(2^{2m}+2^{2m})=2^{3m}$
since there are $2^{2m}$ different $\mathbf{x}$ and $\mathbf{y}$, respectively. This finishes the proof. ∎
6.3. Values for the 33-fold parallel repetition
Now, we will look at the case of 33-fold parallel repetition. We will start with describing a classical strategy with winning probability $\frac{5}{16}$. We note that this strategy performs better that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$. Alice’s and Bob’s strategies are of the form $\left(a_{1},a_{2},a_{3}\right)$, $\left(b_{1},b_{2},b_{3}\right)$ where
$a_{i}:\{0,1\}^{3}\rightarrow\{0,1,\perp\},\text{ }b_{i}:\{0,1\}^{3}\rightarrow\{0,1,\perp\},\quad i=1,2,3$
We will show that the success probability using the following strategy is $5/16$. Given $(x_{1},x_{2},x_{3})$ and $(y_{1},y_{2},y_{3})$, Alice and Bob answer with
$a_{1}$ $=\perp,$ $b_{1}=y_{2}\wedge y_{3},$
$a_{2}$ $=x_{1}\vee x_{3},$ $b_{2}=\perp,$
[…]
respectively. Let us denote the event that Alice and Bob win round $i=1,2,3$ by $W_{i}$. Then
$\Pr[W_{1}]$ $=\Pr[x_{1}=y_{2}\wedge y_{3}]=\frac{1}{2},$
$\Pr[W_{2}|W_{1}]$ $=\Pr[y_{2}=x_{1}\vee x_{3}|x_{1}=y_{2}\wedge y_{3}]=\frac{5}{8}.$
To see the latter, note that
$\Pr[y_{2}=x_{1}\vee x_{3}|x_{1}=y_{2}\wedge y_{3}]$ $=\Pr[y_{2}=(y_{2}\wedge y_{3})\vee x_{3}]$
$=\frac{1}{2}\Pr[y_{3}\vee x_{3}=1]+\frac{1}{2}\Pr[x_{3}=0]$
$=\frac{1}{2}\cdot\frac{3}{4}+\frac{1}{4}$
$=\frac{5}{8},$
where we used that the questions are sampled independently. Now, since
$\Pr[W_{1},W_{2},W_{3}]$ $=\Pr[W_{2}|W_{1},W_{3}]\Pr[W_{1}|W_{3}]\Pr[W_{3}]$
$=\frac{5}{16}\Pr[W_{2}|W_{1},W_{3}],$ (20)
the goal is to show that $\Pr[W_{2}|W_{1},W_{3}]=1$. Assume
$x_{1}=y_{2}\wedge y_{3},$ $y_{2}=x_{1}\vee x_{3}.$
If $x_{1}=y_{2}\wedge y_{3}=0$, the players win round 33 if and only if $y_{2}=x_{3}$. But this is true because $y_{2}=x_{1}\vee x_{3}=x_{3}$ for $x_{1}=0$. If $x_{1}=y_{2}\wedge y_{3}=1$, the players win if and only if $y_{3}=1$, which is true since $y_{2}\wedge y_{3}=1$ implies $y_{3}=1$. This shows $\Pr[W_{2}|W_{1},W_{3}]=1$ and by Equation 20, the strategy has winning probability $\frac{5}{16}$. By exhaustive search through all deterministic strategies using a computer, we have confirmed that this strategy is indeed optimal.
While we see that the previous classical strategy performs better than the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition, there is no non-signalling strategy that outperforms the one where we use optimal non-signalling strategy for one game and the optimal classical strategy for the 22-fold parallel repetition. Therefore, the non-signalling value of the 33-fold parallel repetition is $\frac{1}{3}$. The following is an optimal strategy achieving this winning probability.
[…]
We also obtain $\omega_{ns}(G^{\times 3})=\frac{1}{3}$, which can be verified in a similar fashion as done for the cases $n=1$ and $n=2m$ in this article.
7. Feige’s game as synchronous game
A nonlocal game is called synchronous if the question and answer sets of the players coincide, and additionally they have to return the same answer if they receive the same question in order to win. More formally, it holds $X=Y$, $A=B$ and V(a,b|x,x)=0V(a,b\,|\,x,x)=0 for $a\neq b$. There is the following synchronous version of the Feige game, which we denote by $G_{F}^{s}$:
$X=Y=\{0,1\},\quad A=B=\{A0,A1,B0,B1\},\quad\pi(x,y)=\tfrac{1}{4}$
$V(a,b,x,y)=\begin{cases}1\quad(a,b)=(Ax,Ax)\text{ or }(a,b)=(By,By)\\ 0\quad\text{otherwise}.\end{cases}$
Note that Alice should never answer $A1$, when she receives 0, since the players always lose in this case. The same is true for Bob, which shows that to achieve maximum winning probability, the players will restrict to three answers as in the original Feige game. The difference is that $G_{F}^{s}$ is a synchronous game, whereas $G_{F}$ is not synchronous.
We define a strategy $S$ to be synchronous if p(a,b|x,x)=0p(a,b\,|\,x,x)=0 for $a\neq b$, that is, the players never give different answers when receiving the same question. We can now define the synchronous value $\omega_{q}^{s}(G)$ of a nonlocal game $G$ to be the supremum over all synchronous quantum strategies, i.e.
$\omega_{q}^{s}(G)=\mathrm{sup}\{\omega(S,G)\,|\,S\text{ synchronous quantum strategy}\}.$
The goal of this section is to prove that
$\omega_{q}^{s}(G_{F}^{s})=\frac{1}{2}<\frac{9}{16}=\omega_{q}(G_{F}^{s}).$
Therefore, we have a synchronous game for which the synchronous value is equal to the classical value, but the game has quantum advantage. Such games are known from HMN+ (24), where they considered certain graph coloring games. Note that $\omega_{q}(G_{F}^{s})\geq\frac{9}{16}$ by using the optimal quantum strategy for the Feige game for the synchronous version, and $\omega_{q}(G_{F}^{s})\leq\frac{9}{16}$ follows from the NPA hierarchy at level $1+AB$, in the same way as for $G_{F}$. In the following, we will show $\omega_{q}^{s}(G_{F}^{s})=\frac{1}{2}$. Recall the following lemma from (MPS, 24, Corollary 3.6 a)).
Lemma 7.1.
Let $S=(\mathinner{|\psi\rangle},\{P^{x}_{a}\},\{Q^{y}_{b}\})$ be a synchronous quantum strategy. Then
$P^{x}_{a}\otimes\mathrm{Id}\mathinner{|\psi\rangle}=\mathrm{Id}\otimes Q^{x}_{a}\mathinner{|\psi\rangle}$
for any $x,a$.
With this we can prove our next result.
Proposition 7.2.
It holds $\omega_{q}^{s}(G_{F}^{s})=\frac{1}{2}$.
Proof.
Consider a synchronous quantum strategy $S=(\mathinner{|\psi\rangle},\{P^{x}_{a}\},\{Q^{y}_{b}\})$ for $G_{F}^{s}$. The winning probability of $S$ is given by
$\omega_{q}(S,G_{F}^{s})=\frac{1}{4}\sum_{x,y\in\{0,1\}}\mathinner{\langle\psi|}P^{x}_{Ax}\otimes Q^{y}_{Ax}+P^{x}_{By}\otimes Q^{y}_{By}\mathinner{|\psi\rangle}.$
Since $\mathrm{Id}-P^{x}_{Ax}$ and $\mathrm{Id}-Q^{y}_{By}$ are positive operators, we get $\mathinner{\langle\psi|}P^{x}_{Ax}\otimes Q^{y}_{Ax}\mathinner{|\psi\rangle}\leq\mathinner{\langle\psi|}\mathrm{Id}\otimes Q^{y}_{Ax}\mathinner{|\psi\rangle}$ and $\mathinner{\langle\psi|}P^{x}_{By}\otimes Q^{y}_{By}\mathinner{|\psi\rangle}\leq\mathinner{\langle\psi|}P^{x}_{By}\otimes\mathrm{Id}\mathinner{|\psi\rangle}$. Therefore
$\omega_{q}(S,G_{F}^{s})$ $\leq\frac{1}{4}\sum_{x,y\in\{0,1\}}\mathinner{\langle\psi|}\mathrm{Id}\otimes Q^{y}_{Ax}+P^{x}_{By}\otimes\mathrm{Id}\mathinner{|\psi\rangle}$
$=\frac{1}{4}\sum_{x,y\in\{0,1\}}\mathinner{\langle\psi|}(P^{y}_{Ax}+P^{x}_{By})\otimes\mathrm{Id}\mathinner{|\psi\rangle},$
where we used 7.1 with $Q^{y}_{Ax}$ in the second step. We conclude
$\omega_{q}(S,G_{F}^{s})$ $\leq\frac{1}{4}\mathinner{\langle\psi|}(P^{0}_{A0}+P^{0}_{B0}+P^{0}_{A1}+P^{0}_{B1})+(P^{1}_{A0}+P^{1}_{B0}+P^{1}_{A1}+P^{1}_{B1})\otimes\mathrm{Id}\mathinner{|\psi\rangle}=\frac{1}{2}.$
Acknowledgments
We thank Laura Mančinska for helpful discussions about ”or”-games. We acknowledge support by the European Research Council (ERC Grant SYMOPTIC, 101040907), the German Federal Ministry of Research, Technology and Space (QuSol, 13N17173), and the German Research Foundation under Germany’s Excellence Strategy (EXC 2092 CASA, 390781972). Y.Z. is supported by VILLUM FONDEN via QMATH Centre of Excellence grant number 10059 and Villum Young Investigator grant number 37532.
Appendix A Algebraic structure of Feige’s game from sums of square decompositions
The vector of polynomials
$F=F(p_{a}^{x},q_{b}^{y})=\left(f_{1}(p_{a}^{x},q_{b}^{y}),...,f_{32}(p_{a}^{x},q_{b}^{y})\right)^{T}$
in the proof of 4.4 is given by
$f_{1}(p_{a}^{x},q_{b}^{y})=p_{0}^{0}p_{1}^{1}p_{0}^{0}-\frac{1}{16}p_{0}^{0}$
$f_{2}(p_{a}^{x},q_{b}^{y})=p_{1}^{0}p_{1}^{1}p_{1}^{0}-\frac{9}{16}p_{1}^{0}$
$f_{3}(p_{a}^{x},q_{b}^{y})=p_{0}^{1}p_{1}^{0}p_{0}^{1}-\frac{1}{16}p_{0}^{1}$
$f_{4}(p_{a}^{x},q_{b}^{y})=p_{1}^{1}p_{1}^{0}p_{1}^{1}-\frac{9}{16}p_{1}^{1}$
$f_{5}(p_{a}^{x},q_{b}^{y})=p_{0}^{0}p_{0}^{1}p_{0}^{0}-\frac{9}{16}p_{0}^{0}$
$f_{6}(p_{a}^{x},q_{b}^{y})=p_{0}^{1}p_{0}^{0}p_{0}^{1}-\frac{9}{16}p_{0}^{1}$
$f_{7}(p_{a}^{x},q_{b}^{y})=p_{1}^{0}p_{0}^{1}p_{1}^{0}-\frac{1}{16}p_{1}^{0}$
$f_{8}(p_{a}^{x},q_{b}^{y})=p_{1}^{1}p_{0}^{0}p_{1}^{1}-\frac{1}{16}p_{1}^{1}$
$f_{9}(p_{a}^{x},q_{b}^{y})=\frac{3}{4}p_{0}^{1}p_{1}^{0}+\frac{3}{4}p_{1}^{0}p_{0}^{1}+\frac{5}{4}p_{1}^{0}p_{1}^{1}+\frac{5}{4}p_{1}^{1}p_{1}^{0}+\frac{5}{16}p_{0}^{0}-\frac{3}{8}p_{0}^{1}-\frac{9}{16}p_{1}^{0}-\frac{15}{8}p_{1}^{1}+q_{0}^{0}$
$f_{10}(p_{a}^{x},q_{b}^{y})=\frac{3}{4}p_{0}^{1}p_{1}^{0}+\frac{3}{4}p_{1}^{0}p_{0}^{1}+\frac{5}{4}p_{1}^{0}p_{1}^{1}+\frac{5}{4}p_{1}^{1}p_{1}^{0}-\frac{3}{16}p_{0}^{0}+\frac{1}{8}p_{0}^{1}-\frac{33}{16}p_{1}^{0}-\frac{3}{8}p_{1}^{1}+q_{1}^{0}$
$f_{11}(p_{a}^{x},q_{b}^{y})=\frac{3}{4}p_{0}^{1}p_{1}^{0}+\frac{3}{4}p_{1}^{0}p_{0}^{1}+\frac{5}{4}p_{1}^{0}p_{1}^{1}+\frac{5}{4}p_{1}^{1}p_{1}^{0}+\frac{9}{16}p_{0}^{0}-\frac{9}{8}p_{0}^{1}-\frac{13}{16}p_{1}^{0}-\frac{9}{8}p_{1}^{1}+q_{0}^{1}$
$f_{12}(p_{a}^{x},q_{b}^{y})=\frac{3}{4}p_{0}^{1}p_{1}^{0}+\frac{3}{4}p_{1}^{0}p_{0}^{1}+\frac{5}{4}p_{1}^{0}p_{1}^{1}+\frac{5}{4}p_{1}^{1}p_{1}^{0}-\frac{15}{16}p_{0}^{0}+\frac{3}{8}p_{0}^{1}-\frac{21}{16}p_{1}^{0}-\frac{5}{8}p_{1}^{1}+q_{1}^{1}$
$f_{13}(p_{a}^{x},q_{b}^{y})=I+\frac{2}{3}p_{0}^{1}p_{1}^{0}+\frac{2}{3}p_{1}^{0}p_{0}^{1}+2p_{1}^{0}p_{1}^{1}+2p_{1}^{1}p_{1}^{0}-\frac{1}{2}p_{0}^{0}-\frac{2}{3}p_{0}^{1}-\frac{13}{6}p_{1}^{0}-2p_{1}^{1}$
$f_{14}(p_{a}^{x},q_{b}^{y})=p_{0}^{0}q_{0}^{0}-\frac{3}{8}p_{0}^{1}p_{1}^{0}-\frac{3}{4}p_{1}^{0}p_{0}^{1}+\frac{3}{4}p_{1}^{1}p_{0}^{0}+\frac{1}{8}p_{1}^{1}p_{1}^{0}-\frac{5}{32}p_{0}^{0}+\frac{3}{16}p_{0}^{1}+\frac{3}{32}p_{1}^{0}-\frac{3}{16}p_{1}^{1}$
$f_{15}(p_{a}^{x},q_{b}^{y})=p_{0}^{0}q_{0}^{1}+\frac{3}{8}p_{0}^{1}p_{1}^{0}+\frac{9}{4}p_{1}^{0}p_{0}^{1}-\frac{9}{4}p_{1}^{1}p_{0}^{0}-\frac{5}{8}p_{1}^{1}p_{1}^{0}+\frac{9}{32}p_{0}^{0}-\frac{9}{16}p_{0}^{1}-\frac{3}{32}p_{1}^{0}+\frac{9}{16}p_{1}^{1}$
$f_{16}(p_{a}^{x},q_{b}^{y})=p_{0}^{1}q_{0}^{0}+\frac{9}{8}p_{0}^{1}p_{1}^{0}+\frac{15}{8}p_{1}^{0}p_{0}^{1}+\frac{5}{8}p_{1}^{0}p_{1}^{1}-\frac{15}{8}p_{1}^{1}p_{0}^{0}+\frac{15p_{0}^{0}}{32}-\frac{9p_{0}^{1}}{16}-\frac{15p_{1}^{0}}{32}$
$f_{17}(p_{a}^{x},q_{b}^{y})=p_{0}^{1}q_{0}^{1}+\frac{7}{8}p_{0}^{1}p_{1}^{0}+\frac{21}{8}p_{1}^{0}p_{0}^{1}+\frac{7}{8}p_{1}^{0}p_{1}^{1}-\frac{21}{8}p_{1}^{1}p_{0}^{0}+\frac{21}{32}p_{0}^{0}-\frac{21}{16}p_{0}^{1}-\frac{21}{32}p_{1}^{0}$
$f_{18}(p_{a}^{x},q_{b}^{y})=p_{0}^{0}q_{1}^{0}+\frac{9}{8}p_{0}^{1}p_{1}^{0}-\frac{3}{4}p_{1}^{0}p_{0}^{1}+\frac{3}{4}p_{1}^{1}p_{0}^{0}+\frac{5}{8}p_{1}^{1}p_{1}^{0}-\frac{9}{32}p_{0}^{0}+\frac{3}{16}p_{0}^{1}-\frac{9}{32}p_{1}^{0}-\frac{3}{16}p_{1}^{1}$
$f_{19}(p_{a}^{x},q_{b}^{y})=p_{0}^{0}q_{1}^{1}+\frac{7}{8}p_{0}^{1}p_{1}^{0}-\frac{7}{4}p_{1}^{0}p_{0}^{1}+\frac{7}{4}p_{1}^{1}p_{0}^{0}+\frac{7}{8}p_{1}^{1}p_{1}^{0}-\frac{35}{32}p_{0}^{0}+\frac{7}{16}p_{0}^{1}-\frac{7}{32}p_{1}^{0}-\frac{7}{16}p_{1}^{1}$
$f_{20}(p_{a}^{x},q_{b}^{y})=p_{0}^{1}q_{1}^{0}-\frac{3}{8}p_{0}^{1}p_{1}^{0}+\frac{3}{8}p_{1}^{0}p_{0}^{1}+\frac{1}{8}p_{1}^{0}p_{1}^{1}-\frac{3}{8}p_{1}^{1}p_{0}^{0}+\frac{3}{32}p_{0}^{0}-\frac{1}{16}p_{0}^{1}-\frac{3}{32}p_{1}^{0}$
$f_{21}(p_{a}^{x},q_{b}^{y})=p_{0}^{1}q_{1}^{1}+\frac{3}{8}p_{0}^{1}p_{1}^{0}-\frac{15}{8}p_{1}^{0}p_{0}^{1}-\frac{5}{8}p_{1}^{0}p_{1}^{1}+\frac{15}{8}p_{1}^{1}p_{0}^{0}-\frac{15}{32}p_{0}^{0}+\frac{3}{16}p_{0}^{1}+\frac{15}{32}p_{1}^{0}$
$f_{22}(p_{a}^{x},q_{b}^{y})=p_{1}^{0}q_{0}^{0}+\frac{3}{8}p_{1}^{0}p_{0}^{1}-\frac{5}{8}p_{1}^{0}p_{1}^{1}+\frac{3}{16}p_{1}^{0}$
$f_{23}(p_{a}^{x},q_{b}^{y})=p_{1}^{0}q_{0}^{1}-\frac{3}{8}p_{1}^{0}p_{0}^{1}+\frac{1}{8}p_{1}^{0}p_{1}^{1}-\frac{1}{16}p_{1}^{0}$
$f_{24}(p_{a}^{x},q_{b}^{y})=p_{1}^{1}q_{0}^{0}+\frac{7}{8}p_{1}^{1}p_{0}^{0}+\frac{7}{8}p_{1}^{1}p_{1}^{0}-\frac{21}{16}p_{1}^{1}$
$f_{25}(p_{a}^{x},q_{b}^{y})=p_{1}^{1}q_{0}^{1}+\frac{9}{8}p_{1}^{1}p_{0}^{0}+\frac{5}{8}p_{1}^{1}p_{1}^{0}-\frac{9}{16}p_{1}^{1}$
$f_{26}(p_{a}^{x},q_{b}^{y})=p_{1}^{0}q_{1}^{0}+\frac{7}{8}p_{1}^{0}p_{0}^{1}+\frac{7}{8}p_{1}^{0}p_{1}^{1}-\frac{21}{16}p_{1}^{0}$
$f_{27}(p_{a}^{x},q_{b}^{y})=p_{1}^{0}q_{1}^{1}+\frac{9}{8}p_{1}^{0}p_{0}^{1}+\frac{5}{8}p_{1}^{0}p_{1}^{1}-\frac{9}{16}p_{1}^{0}$
$f_{28}(p_{a}^{x},q_{b}^{y})=p_{1}^{1}q_{1}^{0}+\frac{3}{8}p_{1}^{1}p_{0}^{0}-\frac{5}{8}p_{1}^{1}p_{1}^{0}+\frac{3}{16}p_{1}^{1}$
$f_{29}(p_{a}^{x},q_{b}^{y})=p_{1}^{1}q_{1}^{1}-\frac{3}{8}p_{1}^{1}p_{0}^{0}+\frac{1}{8}p_{1}^{1}p_{1}^{0}-\frac{1}{16}p_{1}^{1}$
$f_{30}(p_{a}^{x},q_{b}^{y})=p_{0}^{0}p_{0}^{1}+3p_{1}^{0}p_{0}^{1}-3p_{1}^{1}p_{0}^{0}-p_{1}^{1}p_{1}^{0}-\frac{3}{4}p_{0}^{1}+\frac{3}{4}p_{1}^{1}$
$f_{31}(p_{a}^{x},q_{b}^{y})=p_{0}^{1}p_{0}^{0}-3p_{1}^{0}p_{0}^{1}-p_{1}^{0}p_{1}^{1}+3p_{1}^{1}p_{0}^{0}-\frac{3}{4}p_{0}^{0}+\frac{3}{4}p_{1}^{0}$
$f_{32}(p_{a}^{x},q_{b}^{y})=p_{0}^{0}p_{1}^{1}-p_{0}^{1}p_{1}^{0}-p_{1}^{0}p_{0}^{1}+p_{1}^{1}p_{0}^{0}-\frac{1}{4}p_{0}^{0}+\frac{1}{4}p_{0}^{1}+\frac{1}{4}p_{1}^{0}-\frac{1}{4}p_{1}^{1}$
Appendix B A generalization of mutually unbiased measurements
For $n\in\mathbb{N}$ and $\nu=(\nu_{ij})_{i,j=1}^{n}$ a real symmetric $n\times n$ matrix with $\nu_{ij}>0$ satisfying $\sum_{j}\nu_{ij}=1$, let $\mathcal{A}_{\nu}$ be the universal $C^{*}$-algebra generated by $E_{1},...,E_{n},F_{1},...,F_{n}$ satisfying:
  • i)
    $E_{i}=E_{i}^{*}$, $F_{i}=F_{i}^{*}$ for all i=1,,ni=1,...,n
  • ii)
    $\sum_{i=1}^{n}E_{i}=\sum_{i=1}^{n}F_{i}=I$
  • iii)
    $E_{i}F_{j}E_{i}=\nu_{ij}E_{i}$, $F_{i}E_{j}F_{i}=\nu_{ij}F_{i}$ for all i,j=1,,ni,j=1,...,n
Observe that by summing over $j$ in iii) and using i) and ii), one easily sees that the generators consist of two sets of orthogonal projections.
Definition B.1.
We call a representation $\rho_{\nu}:\mathcal{A}_{\nu}\rightarrow B(H)$ a $\nu$-biased measurement.
This definition becomes clear when considering the following example
Example B.2.
A pair of $n$-outcome mutually unbiased measurements (MUMs) as defined in TFR+ (21) constitute a representation $\mathcal{A}_{d^{-1}1_{n}}$ where $1_{n}$ is the $n\times n$ matrix of all ones.
Hence $\nu$-biased bases generalize the concept of MUMs (see e.g. TFR+ (21)). It turns out we can generalize results from TFR+ (21) as follows.
Proposition B.3.
Let $\rho_{\nu}:\mathcal{A}_{\nu}\rightarrow B(H)$ ($H$ separable) be a $\nu$-biased measurement, $P_{a}:=\rho(E_{a})$ and $Q_{b}:=\rho(F_{b})$. Then
$H\cong H^{\prime}\otimes\mathbb{C}^{n}$
for some Hilbert space $H^{\prime}$ and
$P_{a}$ $=I_{H^{\prime}}\otimes|a\rangle\langle a|$ (21)
$Q_{b}$ $=\sum_{jk}V_{j}^{(b)}(V_{k}^{(b)})^{*}\otimes|j\rangle\langle k|$ (22)
where the operators $V^{(b)}_{j}$ satisfy
$V_{j}^{(b)}(V_{j}^{(b)})^{*}$ $=\nu_{jb}I_{H^{\prime}}$ (23)
$V_{1}^{(b)}$ $=\sqrt{\nu_{1b}}I_{H^{\prime}}$ (24)
$V_{j}^{(1)}$ $=\sqrt{\nu_{1j}}I_{H^{\prime}}$ (25)
$\sum_{k}(V_{k}^{(b)})^{*}V_{k}^{(b^{\prime})}$ $=0,\quad\forall b\neq b$ (26)
$\sum_{b}V_{j}^{(b)}(V_{k}^{(b)})^{*}$ $=\delta_{jk}I_{H^{\prime}}$ (27)
Proof.
The operator $T_{ab}:=\tfrac{1}{\sqrt{\nu_{ab}}}P_{a}Q_{b}$ is a partial isometry with support projection $Q_{b}$ and range projection $P_{a}$ i. e.
$T_{ab}T_{ab}^{*}=P_{a},\quad T_{ab}^{*}T_{ab}=Q_{b}.$
This implies that either all projections are finite rank and $\operatorname{Tr}(P_{i})=\operatorname{Tr}(Q_{j})=:\tau$ for all i,j=1,,ni,j=1,...,n or none of them are. Notice that for finite dimensional representations we immediately have that the dimension is a multiple of $n$. Indeed,
$n\tau=\sum_{i=1}^{n}\operatorname{Tr}\left(P_{i}\right)=\operatorname{Tr}(I)=d.$
The completeness relation $\sum_{i}P_{i}=I$ and the above considerations allow us to write
$H\cong\bigoplus_{i=1}^{n}H_{i}\cong H^{\prime}\otimes\mathbb{C}^{n}$
where $H_{i}\subseteq H$ is the subspace, $P_{i}$ projects onto and $H^{\prime}$ is another (potentially infinite dimensional) Hilbert space. Now we write
$P_{a}=I_{H^{\prime}}\otimes|a\rangle\langle a|,\quad Q_{b}=\sum_{ij}X_{ij}^{(b)}\otimes|i\rangle\langle j|$
for some ONB $\{\mathinner{|a\rangle}\}$ and some operators $X_{ij}^{(b)}\in B(H^{\prime})$. First note that
$(X_{ij}^{(b)})^{*}=X_{ji}^{(b)}$
since $Q_{b}$ is self-adjoint. Now, by assumption,
$P_{a}Q_{b}P_{a}=X_{aa}^{(b)}\otimes|a\rangle\langle a|=\nu_{ab}I_{H^{\prime}}\otimes|a\rangle\langle a|$
which implies
$X_{aa}^{(b)}=\nu_{ab}I_{H^{\prime}}$ (28)
and
$Q_{b}P_{a}Q_{b}=\sum_{ij}X_{ia}^{(b)}X_{aj}^{(b)}\otimes|i\rangle\langle j|=\sum_{ij}\nu_{ab}X_{ij}^{(b)}\otimes|i\rangle\langle j|$
which implies
$X_{ia}^{(b)}X_{aj}^{(b)}=\nu_{ab}X_{ij}^{(b)}$ (29)
We will use (28)-(29) several times in the following. We claim now that the operator,
$U:=\tfrac{1}{\sqrt{\nu_{11}}}\sum_{j}\tfrac{1}{\sqrt{\nu_{1j}}}X_{1j}^{(1)}\otimes|j\rangle\langle j|$
is unitary. Indeed,
$UU^{*}=\tfrac{1}{\nu_{11}}\sum_{j}\tfrac{1}{\nu_{1j}}X_{1j}^{(1)}X_{j1}^{(1)}\otimes|j\rangle\langle j|=\tfrac{1}{\nu_{11}}\sum_{j}X_{11}^{(1)}\otimes|j\rangle\langle j|=\sum_{j}I_{H^{\prime}}\otimes|j\rangle\langle j|=I_{H}$
and similarly $U^{*}U=I_{H}$. Then we have
$UP_{a}U^{*}=\tfrac{1}{\nu_{11}}\tfrac{1}{\nu_{1a}}X_{1a}^{(1)}X_{a1}^{(1)}\otimes|j\rangle\langle j|=\tfrac{1}{\nu_{11}}X_{11}^{(1)}\otimes|j\rangle\langle j|=P_{a}$
and
$Q_{b}^{\prime}:=UQ_{b}U^{*}=\tfrac{1}{\nu_{11}}\sum_{jk}\tfrac{1}{\sqrt{\nu_{1j}\nu_{1k}}}X_{1j}^{(1)}X_{jk}^{(b)}X_{k1}^{(1)}\otimes|j\rangle\langle k|$
In particular we find
$\begin{split}Q_{1}^{\prime}=&\tfrac{1}{\nu_{11}}\sum_{jk}\tfrac{1}{\sqrt{\nu_{1j}\nu_{1k}}}X_{1j}^{(1)}X_{jk}^{(1)}X_{k1}^{(1)}\otimes|j\rangle\langle k|=\tfrac{1}{\nu_{11}}\sum_{jk}\tfrac{\sqrt{\nu_{1j}}}{\sqrt{\nu_{1k}}}X_{1k}^{(1)}X_{k1}^{(1)}\otimes|j\rangle\langle k|\\ =&\tfrac{1}{\nu_{11}}\sum_{jk}\sqrt{\nu_{1j}}\sqrt{\nu_{1k}}X_{11}^{(1)}\otimes|j\rangle\langle k|=\sum_{jk}\sqrt{\nu_{1j}}\sqrt{\nu_{1k}}I_{H^{\prime}}\otimes|j\rangle\langle k|=I_{H^{\prime}}\otimes|v\rangle\langle v|\end{split}$
where $\mathinner{|v\rangle}:=\sum_{j}\sqrt{\nu_{1j}}\mathinner{|j\rangle}$ is a unit vector. Now we characterize the remaining $Q_{b}^{\prime}$. Let
$Y_{jk}^{(b)}=\tfrac{1}{\nu_{11}\sqrt{\nu_{1j}\nu_{1k}}}X_{1j}^{(1)}X_{jk}^{(b)}X_{k1}^{(1)}$
so that
$Q_{b}^{\prime}=\sum_{jk}Y_{jk}^{(b)}\otimes|j\rangle\langle k|$
and notice that
$\begin{split}Y_{ja}^{(b)}Y_{ak}^{(b)}=&\tfrac{1}{\nu_{11}^{2}\nu_{1a}\sqrt{\nu_{1j}\nu_{1k}}}(X_{1j}^{(1)}X_{ja}^{(b)}X_{a1}^{(1)})(X_{1a}^{(1)}X_{ak}^{(b)}X_{k1}^{(1)})=\tfrac{1}{\nu_{11}\sqrt{\nu_{1j}\nu_{1k}}}X_{1j}^{(1)}X_{ja}^{(b)}X_{ak}^{(b)}X_{k1}^{(1)}\\ =&\tfrac{\nu_{ab}}{\nu_{11}\sqrt{\nu_{1j}\nu_{1k}}}X_{1j}^{(1)}X_{jk}^{(b)}X_{k1}^{(1)}=\nu_{ab}Y_{jk}^{(b)}\end{split}$ (30)
Hence we can write
$Q_{b}^{\prime}=\sum_{jk}V_{j}^{(b)}(V_{k}^{(b)})^{*}\otimes|j\rangle\langle k|$
where
$V_{j}^{(b)}:=\tfrac{1}{\sqrt{\nu_{1b}}}Y_{j1}^{(b)}=\tfrac{1}{\sqrt{\nu_{1b}\nu_{1j}\nu_{11}}}X_{1j}^{(1)}X_{j1}^{(b)}.$
so
$V_{1}^{(b)}=\tfrac{1}{\sqrt{\nu_{1b}\nu_{11}\nu_{11}}}X_{11}^{(1)}X_{11}^{(b)}=\sqrt{\nu_{1b}}I$
and
$V_{j}^{(1)}:=\tfrac{1}{\sqrt{\nu_{11}}}Y_{j1}^{(1)}=\tfrac{1}{\sqrt{\nu_{11}\nu_{1j}\nu_{11}}}X_{1j}^{(1)}X_{j1}^{(1)}=\sqrt{\nu_{1j}}I.$
Moreover, we have
$\begin{split}V_{j}^{(b)}(V_{j}^{(b)})^{*}=&\tfrac{1}{\nu_{1b}\nu_{1j}\nu_{11}}X_{1j}^{(1)}X_{j1}^{(b)}X_{1j}^{(b)}X_{j1}^{(1)}=\tfrac{1}{\nu_{1j}\nu_{11}}X_{1j}^{(1)}X_{jj}^{(b)}X_{j1}^{(1)}\\ =&\tfrac{\nu_{jb}}{\nu_{1j}\nu_{11}}X_{1j}^{(1)}X_{j1}^{(1)}=\tfrac{\nu_{jb}}{\nu_{11}}X_{11}^{(1)}=\nu_{jb}I\end{split}$
Orthogonality of the $Q_{b}^{\prime}$’s gives us
$\begin{split}Q_{b}^{\prime}Q_{b^{\prime}}^{\prime}&=\sum_{jk^{\prime}}V_{j}^{(b)}\left[\sum_{k}(V_{k}^{(b)})^{*}V_{k}^{(b^{\prime})}\right](V_{k^{\prime}}^{(b^{\prime})})^{*}\otimes|j\rangle\langle k^{\prime}|\\ &=V_{1}^{(b)}\left[\sum_{k}(V_{k}^{(b)})^{*}V_{k}^{(b^{\prime})}\right](V_{1}^{(b^{\prime})})^{*}\otimes|1\rangle\langle 1|+...\\ &=\sqrt{\nu_{1b}\nu_{1b^{\prime}}}\left[\sum_{k}(V_{k}^{(b)})^{*}V_{k}^{(b^{\prime})}\right]\otimes|1\rangle\langle 1|+...\\ &=0\end{split}$
which implies
$\sum_{k}(V_{k}^{(b)})^{*}V_{k}^{(b^{\prime})}=0,\quad\forall b\neq b^{\prime}$
Completeness gives us
$\begin{split}I_{H}=\sum_{jk}\left[\sum_{b}V_{j}^{(b)}(V_{k}^{(b)})^{*}\right]\otimes|j\rangle\langle k|=\sum_{j}I_{H^{\prime}}\otimes|j\rangle\langle j|\end{split}$
and hence
$\sum_{b}V_{j}^{(b)}(V_{k}^{(b)})^{*}=\delta_{jk}I_{H^{\prime}}$
Proposition B.4.
Let $\rho:\mathcal{A}_{\nu}\rightarrow B(H)$ ($H$ separable) be a representation, $P_{a}:=\rho(E_{a})$ and $Q_{b}:=\rho(F_{b})$. If the $V_{j}^{(b)}$’s commute, then the $P_{a}$’s and $Q_{b}$’s are direct sums of rank 1 projections.
Proof.
We have
$Y_{jk}^{(b)},Y_{j^{\prime}k^{\prime}}^{(b^{\prime})}]=0$
So there’s an ONB $\mathinner{|e_{j}\rangle}$ for $H^{\prime}$ in which
$Y_{jk}^{(b)}=\sum_{\ell}(\lambda_{jk}^{(b)})_{\ell}|e_{\ell}\rangle\langle e_{\ell}|$
and
$Q_{b}^{\prime}=\sum_{jk}\sum_{\ell}(\lambda_{jk}^{(b)})_{\ell}|e_{\ell}\rangle\langle e_{\ell}|\otimes|j\rangle\langle k|=\sum_{\ell}|e_{\ell}\rangle\langle e_{\ell}|\otimes T_{\ell}^{(b)}$
where
$T_{\ell}^{(b)}:=\sum_{jk}(\lambda_{jk}^{(b)})_{\ell}|j\rangle\langle k|$
Note
$Y_{ja}^{(b)}Y_{ak}^{(b)}=\sum_{\ell}(\lambda_{ja}^{(b)})_{\ell}(\lambda_{ak}^{(b)})_{\ell}|e_{\ell}\rangle\langle e_{\ell}|=\sum_{\ell}\nu_{ab}(\lambda_{jk}^{(b)})_{\ell}|e_{\ell}\rangle\langle e_{\ell}|$
so in particular we have
$(\lambda_{ja}^{(b)})_{\ell}(\lambda_{ak}^{(b)})_{\ell}=\nu_{ab}(\lambda_{jk}^{(b)})_{\ell}$
Since $Q_{b}=Q_{b}^{*}=Q_{b}^{2}$ we must have $T_{\ell}^{(b)}=(T_{\ell}^{(b)})^{*}=(T_{\ell}^{(b)})^{2}$. We have from earlier that
$Y_{jj}^{(b)}=\tfrac{\nu_{jb}}{\nu_{11}\nu_{1j}}X_{1j}^{(1)}X_{jj}^{(b)}X_{j1}^{(1)}=\tfrac{\nu_{jb}}{\nu_{11}\nu_{1j}}X_{1j}^{(1)}X_{j1}^{(1)}=\nu_{jb}I_{H^{\prime}}$
hence we must have $(\lambda_{jj}^{(b)})_{\ell}=\nu_{jb}$ and so
$\operatorname{Tr}(T_{\ell}^{(b)})=\sum_{j}(\lambda_{jj}^{(b)})_{\ell}=\sum_{j}\nu_{jb}=1$
so the $T_{\ell}^{(b)}$’s are rank one projections. Now, notice that
$|a\rangle\langle a|T_{\ell}^{(b)}|a\rangle\langle a|=\sum_{jk}(\lambda_{jk}^{(b)})_{\ell}\mathinner{\langle a|j\rangle}\mathinner{\langle k|a\rangle}|a\rangle\langle a|=(\lambda_{aa}^{(b)})_{\ell}|a\rangle\langle a|=\nu_{ab}|a\rangle\langle a|$
and
$\begin{split}T_{\ell}^{(b)}|a\rangle\langle a|T_{\ell}^{(b)}=&\sum_{jkj^{\prime}k^{\prime}}(\lambda_{jk}^{(b)})_{\ell}(\lambda_{j^{\prime}k^{\prime}}^{(b)})_{\ell}|j\rangle\langle k||a\rangle\langle a||j^{\prime}\rangle\langle k^{\prime}|\\ =&\sum_{jk^{\prime}}(\lambda_{ja}^{(b)})_{\ell}(\lambda_{ak^{\prime}}^{(b)})_{\ell}|j\rangle\langle k^{\prime}|\\ =&\nu_{ab}\sum_{jk^{\prime}}(\lambda_{jk^{\prime}}^{(b)})_{\ell}|j\rangle\langle k^{\prime}|=\nu_{ab}T_{\ell}^{(b)}\end{split}$
We conclude that $\{|a\rangle\langle a|\}_{a=1}^{n}$ and $\{T_{\ell}^{(b)}\}_{b=1}^{n}$ are rank 1 representations of $\mathcal{A}_{\nu}$ for all $\ell$. ∎
B.1. Equivalences
Let us refer to rank 1 representations of $\mathcal{A}_{\nu}$ as $\nu$-biased bases. Following BBE+ (07) we define the following equivalence relation:
Definition B.5 (Equivalence of $\nu$-biased measurements, $\sim$).
Let
$(\{\mathinner{|\varphi_{i}^{(1)}\rangle}\}_{i=1}^{n},\{\mathinner{|\psi_{j}^{(1)}\rangle}\}_{j=1}^{n})\text{ and }(\{\mathinner{|\varphi_{i}^{(2)}\rangle}\}_{i=1}^{n},\{\mathinner{|\psi_{j}^{(2)}\rangle}\}_{j=1}^{n})$
be two pairs of $n$-dimensional $\nu$-biased bases i. e.
$\lvert\mathinner{\langle\varphi_{i}^{(k)}|\psi_{j}^{(k)}\rangle}\rvert^{2}=\nu_{ij}$
for $k=1,2$. Let
$M_{k}:=(\mathinner{|\varphi_{i}^{(k)}\rangle}_{j})_{ij},\quad N_{k}:=(\mathinner{|\psi_{i}^{(k)}\rangle}_{j})_{ij}.$
Then these pairs of $\nu$-biased bases are equivalent denoted
$(M_{1},N_{1})\sim(M_{2},N_{2})$
if and only if there exist permutation matrices $P,P^{\prime}$, diagonal unitaries $D,D^{\prime}$ and a unitary $U$ such that
$M_{2}=UM_{1}PD,\quad N_{2}=UN_{1}P^{\prime}D^{\prime}.$
We also need the following definition.
Definition B.6 (Equivalence of unitaries, $\approx$).
Two unitary matrices $U_{1}$ and $U_{2}$ are equivalent, written $U_{1}\approx U_{2}$ if and only if there exist permutation matrices $P,P^{\prime}$ and diagonal unitaries $D,D^{\prime}$ such that
$U_{2}=PDU_{1}D^{\prime}P^{\prime}.$
This equivalence reflects the fact that permutation of basis elements as well as multiplication with phase factors leave the basis in question essentially unchanged. Now, it is clear that for any pair of $\nu$-biased bases $(M,N)$ we have
$(M,N)\sim(I,M^{*}N)$
It is also clear that since $M^{*}N$ is a basis, which is $\nu$-biased with the computational basis it must be given by a (generalized) complex Hadamard matrix
$M^{*}N=H=(h_{ij})_{ij}$
where $\lvert h_{ij}\rvert=\sqrt{\nu_{ij}}$ and $HH^{*}=I_{n}$. Hence, a pair of $\nu$-biased bases can always be written in standard form, $(I,H)$, for some (generalized) complex Hadamard matrix. It follows by the definitions that two pairs of $\nu$-biased bases, $(I,H)$ and $(I,H^{\prime})$ are $\sim$-equivalent if and only if $H\approx H^{\prime}$.
In the following, we will study in which cases $\nu$-biased bases are unique up to $\sim$-equivalence.
Lemma B.7.
Let $a_{1},\dots,a_{n}>0$ be strictly positive. If it holds $\sum_{k\neq l}a_{k}=a_{l}$ for some $1\leq l\leq n$, then the equation
$\sum_{j=1}^{n-1}e^{2\pi i\phi_{j}}a_{j}+a_{n}=0$ (31)
has a unique solution for $\phi_{j}\in[0,1)$, $1\leq j\leq n-1$.
Proof.
Assume $\sum_{k\neq l}a_{k}=a_{l}$ holds for some $l$. Multiplying Equation 31 with $e^{-2\pi i\phi_{l}}$, we obtain
$\sum_{j\neq l}e^{2\pi i(\phi_{j}-\phi_{l})}a_{j}+a_{l}=0,$
where we $\phi_{n}=0$. From $\sum_{k\neq l}a_{k}=a_{l}$, we get
$0=\sum_{j\neq l}e^{2\pi i(\phi_{j}-\phi_{l})}a_{j}+a_{l}=\sum_{j\neq l}e^{2\pi i(\phi_{j}-\phi_{l})}a_{j}+\sum_{j\neq l}a_{j}=\sum_{j\neq l}(e^{2\pi i(\phi_{j}-\phi_{l})}+1)a_{j}.$
Since $a_{j}>0$, we obtain $\mathrm{Re}(e^{2\pi i(\phi_{j}-\phi_{l})})=-1$, which yields $e^{2\pi i(\phi_{j}-\phi_{l})}=-1$. Especially, we have $e^{-2\pi i\phi_{l}}=e^{2\pi i(\phi_{n}-\phi_{l})}=-1$. This shows $\phi_{l}=\frac{1}{2}$ and we deduce $\phi_{j}=0$ for $j\neq l$. ∎
It is known for example from Haa (97), that for $n=2,3$, all Hadamard matrices are equivalent. The next proposition gives a generalization of this result.
Proposition B.8.
Consider a matrix $\nu=\{\nu_{ij}\}_{ij}$ with positive real entries such that there is a unitary matrix $H$ with $|h_{ij}|=\sqrt{\nu_{ij}}$. If for every $a,b$, there exists $l$ such that
$\sum_{k\neq l}\sqrt{\nu_{ak}}\sqrt{\nu_{bk}}=\sqrt{\nu_{al}}\sqrt{\nu_{bl}},$ (32)
then $H^{\prime}\approx H$ for every unitary $H^{\prime}$ with $|h^{\prime}_{ij}|=\sqrt{\nu_{ij}}$.
Proof.
Let $\nu$ be as in the statement and consider $H$ with $|h_{ij}|=\sqrt{\nu_{ij}}$. By multiplying with appropriate diagonal matrices $D,D^{\prime}$ from left and right, we get
$DHD^{\prime}=\begin{pmatrix}\sqrt{\nu}_{11}&\sqrt{\nu}_{12}&\dots&\sqrt{\nu}_{1n}\\ \sqrt{\nu}_{21}&e^{2\pi i\phi_{22}}\sqrt{\nu}_{22}&\dots&e^{2\pi i\phi_{2n}}\sqrt{\nu}_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ \sqrt{\nu}_{n1}&e^{2\pi i\phi_{n2}}\sqrt{\nu}_{n2}&\dots&e^{2\pi i\phi_{nn}}\sqrt{\nu}_{nn}.\end{pmatrix}$
Since $DHD^{\prime}$ is a unitary matrix, we have
$\sum_{k=2}^{n}e^{2\pi i\phi_{ak}}\sqrt{\nu_{1k}}\sqrt{\nu_{ak}}+\sqrt{\nu_{11}}\sqrt{\nu_{a1}}=0$
for all $2\leq a\leq n$. Because of Equation 32, we can use B.7 and get that all equations above have unique solutions for $\phi_{ak}\in[0,1)$, $2\leq a,k\leq n$. Therefore $DHD^{\prime}$ is unique. ∎
Theorem B.9.
Consider a matrix $\nu=\{\nu_{ij}\}_{ij}\in M_{n}(\mathbb{C})$ for $n=2,3$ with positive real entries such that there is a unitary matrix $H$ with $|h_{ij}|=\sqrt{\nu_{ij}}$ and Equation 32 holds. Then the algebra $\mathcal{A}_{\nu}$ has a unique irreducible representation up to $\sim$-equivalence, which is of dimension $n$.
Proof.
Due to B.4 it suffices to show that the $V_{j}^{(b)}$’s commute. First note that since
$V_{j}^{(1)}=\sqrt{\nu_{1j}}I_{H^{\prime}}\text{ and }\sum_{k}(V_{k}^{(b)})^{*}V_{k}^{(b^{\prime})}=0,\ \forall b\neq b^{\prime}$
we must have
$\sum_{k}V_{k}^{(b)}=0,\ \forall b\neq 1$
and since
$V_{1}^{(b)}=\sqrt{\nu_{1b}}I_{H^{\prime}}\text{ and }\sum_{b}V_{j}^{(b)}(V_{k}^{(b)})^{*}=\delta_{jk}I_{H^{\prime}}$
we must also have
$\sum_{b}V_{j}^{(b)}=0,\ \forall j\neq 1.$
For $n=2$ we have
$V_{1}^{(1)}=\sqrt{\nu_{11}}I_{H^{\prime}},\ V_{2}^{(1)}=\sqrt{\nu_{12}}I_{H^{\prime}},\ V_{1}^{(2)}=\sqrt{\nu_{12}}I_{H^{\prime}},\ V_{2}^{(2)}=-\sqrt{\nu_{12}}I_{H^{\prime}}.$
For $n=3$ we have
$\begin{split}V_{1}^{(1)}=&\sqrt{\nu_{11}}I_{H^{\prime}}\\ V_{2}^{(1)}=V_{1}^{(2)}=&\sqrt{\nu_{12}}I_{H^{\prime}}\\ V_{3}^{(1)}=V_{1}^{(3)}=&\sqrt{\nu_{13}}I_{H^{\prime}}\\ V_{2}^{(2)}=:&V\\ V_{3}^{(2)}=V_{2}^{(3)}=&-V-\sqrt{\nu_{12}}I_{H^{\prime}}\\ V_{3}^{(3)}=&V+(\sqrt{\nu_{12}}-\sqrt{\nu_{13}})I_{H^{\prime}}\end{split}$
which clearly all commute. Now, $P_{a}$ and $Q_{b}$ are direct sums of rank one $\nu$-biased measurements. The desired now follows from Proposition B.8. ∎
For $\nu=\begin{bmatrix}\frac{9}{16}&\frac{1}{16}&\frac{3}{8}\\ \frac{1}{16}&\frac{9}{16}&\frac{3}{8}\\ \frac{3}{8}&\frac{3}{8}&\frac{1}{4}\end{bmatrix}$, we further prove that $\mathcal{A}_{\nu}$ has a unique irreducible representation up to unitary equivalence.
Proof of 4.5.
[…]
Suppose $\pi$ is another irreducible representation of $\mathcal{A}_{\nu}$. By B.9, $\pi$ is $\sim$ equivalent to $\tilde{\pi}$. This means there is a pair of $\nu$-biased bases $\{\{\mathinner{|e_{i}\rangle}\}_{i=1}^{3},\{\mathinner{|f_{i}\rangle}_{i=1}^{3}\}\}$ such that $\pi(E_{i})=\mathinner{|e_{i}\rangle}\mathinner{\langle e_{i}|}$ and $\pi(F_{i})=\mathinner{|f_{i}\rangle}\mathinner{\langle f_{i}|}$ for all $1\leq i\leq 3$, and that there exist permutations $\sigma,\sigma^{\prime}\in S_{3}$ and a unitary $U$ satisfy
$\tilde{\pi}(E_{i})=U\pi(E_{\sigma(i)})U^{*}\text{ and }\tilde{\pi}(F_{i})=U\pi(F_{\sigma^{\prime}(i)})U^{*}$ (33)
[…]

### v2

@@
 In this article, we study a nonlocal game with two questions and three answers per player, which was first considered by Feige in 1991, and show that there is quantum advantage in this game.
-We prove that the game is a robust self-test for the 33-dimensional maximally entangled state.
+We prove that the game is a robust self-test for the $3$-dimensional maximally entangled state.
 Furthermore, we show that the game can be seen as the ”or” of two games that each do not have quantum advantage.
@@
 Note that the relations look similar to the relations obtained from mutually unbiased measurements TFR+ (21), where for the latter it holds $\nu=\frac{1}{d}I$.
-In case of $P^{x}_{a}=\mathinner{|\phi^{x}_{a}\rangle}\mathinner{\langle\phi_{a}^{x}|}$ being rank 11 projections, the transition matrix from the basis $\{\mathinner{|\phi^{x}_{a}\rangle}\}$ to the basis $\{\mathinner{|\phi^{1-x}_{a}\rangle}\}$ is given by a unitary matrix $H$ with $|h_{aa^{\prime}}|=\sqrt{\nu_{aa^{\prime}}}$ (which is a Hadamard matrix in the case of $\nu=\frac{1}{d}I$).
+In case of $P^{x}_{a}=\mathinner{|\phi^{x}_{a}\rangle}\mathinner{\langle\phi_{a}^{x}|}$ being rank $1$ projections, the transition matrix from the basis $\{\mathinner{|\phi^{x}_{a}\rangle}\}$ to the basis $\{\mathinner{|\phi^{1-x}_{a}\rangle}\}$ is given by a unitary matrix $H$ with $|h_{aa^{\prime}}|=\sqrt{\nu_{aa^{\prime}}}$ (which is a Hadamard matrix in the case of $\nu=\frac{1}{d}I$).
 Similar to some Hadamard matrices being unique up to equivalence Haa (97), we show that the same is true for unitaries with $|h_{aa^{\prime}}|=\sqrt{\nu_{aa^{\prime}}}$ for the specific choice of $\nu$ from above.
@@
 Question.
-Does the quantum value exceed the classical value for the 33-fold parallel repetition of Feige’s game?
-We are not aware of a game with quantum advantage, for which the classical and quantum value agree for the 22-fold parallel repetition, but then they separate again for $n=3$.
+Does the quantum value exceed the classical value for the $3$-fold parallel repetition of Feige’s game?
+We are not aware of a game with quantum advantage, for which the classical and quantum value agree for the $2$-fold parallel repetition, but then they separate again for $n=3$.
 As we can see in Table 1 this happens for the classical and the non-signalling value of Feige’s game.
-Note that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition performs worse than the optimal classical strategy for $n=3$, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$.
-Therefore, there is no natural candidate of a quantum strategy that yields quantum advantage for the 33-fold parallel repetition.
-The $\frac{5}{16}$ optimal classical strategy is also interesting: It performs worse on the first two rounds of the game in comparison to the optimal strategy for the 22-fold parallel repetition, but is constructed such that if the players win the first two rounds, they win the third round automatically.
+Note that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the $2$-fold parallel repetition performs worse than the optimal classical strategy for $n=3$, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$.
+Therefore, there is no natural candidate of a quantum strategy that yields quantum advantage for the $3$-fold parallel repetition.
+The $\frac{5}{16}$ optimal classical strategy is also interesting: It performs worse on the first two rounds of the game in comparison to the optimal strategy for the $2$-fold parallel repetition, but is constructed such that if the players win the first two rounds, they win the third round automatically.
 2. Preliminaries
@@
 There are different types of strategies, which we now describe. We start with classical strategies.
-Definition $2.$
+Definition 2.2.
 A classical strategy for a nonlocal game $G$ consists of
@@
   2. (2)
-    probability distributions $\{p_{\omega}(a|x):x\in X,\omega\in\Omega\}$ with outcomes in AA and $\{q_{\omega}(b|y):y\in Y,\omega\in\Omega\}$ with outcomes in $B$.
+    probability distributions $\{p_{\omega}(a|x):x\in X,\omega\in\Omega\}$ with outcomes in $A$ and $\{q_{\omega}(b|y):y\in Y,\omega\in\Omega\}$ with outcomes in $B$.
 In this case, we have
@@
 The proof of our main theorem follows an operator-algebraic framework for self-testing developed inZha (24).
-For finite sets $X$ and AA, we use $\mathscr{A}_{PVM}^{X,A}$ to denote the universal $C^{*}$-algebra generated by $\{p^{x}_{a}:a\in A,x\in X\}$, subject to the relations
+For finite sets $X$ and $A$, we use $\mathscr{A}_{PVM}^{X,A}$ to denote the universal $C^{*}$-algebra generated by $\{p^{x}_{a}:a\in A,x\in X\}$, subject to the relations
   • $(p^{x}_{a})^{2}=(p^{x}_{a})^{*}=p^{x}_{a}\quad\forall a\in A,x\in X$
@@
     Once a player answers a bit given a question (e.g.
-Alice answers 0 when receiving 0), the game will always be lost if the other player receives a question different from this bit (the referee sends 11 to Bob), which happens halve of the time.
+Alice answers 0 when receiving 0), the game will always be lost if the other player receives a question different from this bit (the referee sends $1$ to Bob), which happens halve of the time.
   • (ii)
@@
 We have stored this exact solution in a Mathematica notebook in a GitHub repository SoS . Running the notebook verifies the claims made in this paper.
-$3.$ The non-signalling value of Feige’s game
+3.3. The non-signalling value of Feige’s game
 We have seen that the classical value of Feige’s game is $\frac{1}{2}$, and we can exceed this value with a quantum strategy that wins with probability $\frac{9}{16}$.
@@
 We get the following algebraic structure of optimal strategies for $G_{F}$.
-Lemma $4.$
+Lemma 4.4.
 Let $\nu:=\begin{bmatrix}\frac{9}{16}&\frac{1}{16}&\frac{3}{8}\\ \frac{1}{16}&\frac{9}{16}&\frac{3}{8}\\ \frac{3}{8}&\frac{3}{8}&\frac{1}{4}\end{bmatrix}$, with the first, second, and third column/row labelled by $0,1$, and $\bot$, respectively.
@@
     $E_{i}F_{j}E_{i}=\nu_{ij}E_{i}$, $F_{i}E_{j}F_{i}=\nu_{ij}F_{i}$ for all $i,j$,
-where $\nu$ is the symmetric stochastic matrix defined in $4.$ In fact, for any $n\times n$ symmetric stochastic matrix $\nu$, we can define $\mathcal{A}_{\nu}$ analogously, where generators $\{E_{i}\}_{i=1}^{n}$ and $\{F_{i}\}_{i=1}^{n}$ are PVMs satisfying $E_{i}F_{j}E_{i}=\nu_{ij}E_{i}$ and $F_{i}E_{j}F_{i}=\nu_{ij}F_{i}$.
+where $\nu$ is the symmetric stochastic matrix defined in 4.4.
+In fact, for any $n\times n$ symmetric stochastic matrix $\nu$, we can define $\mathcal{A}_{\nu}$ analogously, where generators $\{E_{i}\}_{i=1}^{n}$ and $\{F_{i}\}_{i=1}^{n}$ are PVMs satisfying $E_{i}F_{j}E_{i}=\nu_{ij}E_{i}$ and $F_{i}E_{j}F_{i}=\nu_{ij}F_{i}$.
 In Appendix B, we establish the representation theory for those $\mathcal{A}_{\nu}$’s.
@@
 Proof.
-Let $(\Gamma,\mathcal{R})$ be the $O(\sqrt{\epsilon})$-robust determining pair of $G_{F}$ defined in $4.$ By 4.5, the associated game algebra $C^{*}(G)=\mathscr{A}_{PVM}^{\{0,1\},\{0,1,\bot\}}/\langle\mathcal{R}\rangle$ is a full matrix algebra.
+Let $(\Gamma,\mathcal{R})$ be the $O(\sqrt{\epsilon})$-robust determining pair of $G_{F}$ defined in 4.4.
+By 4.5, the associated game algebra $C^{*}(G)=\mathscr{A}_{PVM}^{\{0,1\},\{0,1,\bot\}}/\langle\mathcal{R}\rangle$ is a full matrix algebra.
 It follows from 4.3 that $G_{F}$ is a robust self-test.
@@
 We do not know if there is quantum advantage for $n=3$.
-Note that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition performs worse than the optimal classical strategy for $n=3$, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$.
+Note that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the $2$-fold parallel repetition performs worse than the optimal classical strategy for $n=3$, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$.
 We first recall the definition of parallel repetition. In the following, we define the product of two games, which then leads to the definition of the parallel repetition of a game.
@@
 We see that the players win if and only if $x_{1}=y_{2}$, which happens with probability $\frac{1}{2}$.
-This may be surprising, since the classical winning probability of the 22-fold parallel repetition does not decrease when compared with the original game $G_{F}$.
+This may be surprising, since the classical winning probability of the $2$-fold parallel repetition does not decrease when compared with the original game $G_{F}$.
 For the $2m$-fold parallel repetition $G_{F}^{\times 2m}$, we obtain a classical strategy with winning probability $\frac{1}{2^{m}}=\frac{1}{2^{n/2}}$ by repeating the strategy above $m$ times.
@@
 We define $\mathbf{x}|_{I}\in\{0,1\}^{|I|}$ to be the restriction of $\mathbf{x}$ to a subset $I\subseteq[n]$.
-Furthermore, we write $|I|$ for the cardinality of a set II and denote the complement by $I^{c}=[n]\setminus I$.
+Furthermore, we write $|I|$ for the cardinality of a set $I$ and denote the complement by $I^{c}=[n]\setminus I$.
 Lemma 6.3.
@@
 $\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{z},\mathbf{y})$ (15)
-because of $p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$ and the non-signalling property of pp.
+because of $p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$ and the non-signalling property of $p$.
 Since $b_{I}(\mathbf{v})=b_{I}(\mathbf{w})$ if and only if $\mathbf{v}|_{I}=\mathbf{w}|_{I}$, we see that the right hand side of Equation 15 is the same for each $\mathbf{x}|_{I^{c}}\in J$.
@@
 Proof.
-Since $n=2m$, both sets II and $I^{c}$ have $m$ elements.
+Since $n=2m$, both sets $I$ and $I^{c}$ have $m$ elements.
 Therefore, we can choose pairs $(i_{k},j_{k})$ for $1\leq k\leq m$ with $i_{k}\in I,j_{k}\in I^{c}$ such that $\{i_{k},j_{k}|1\leq k\leq m\}=[n]$.
@@
 $2^{4m}\omega(G_{F}^{\times n},p)=\sum_{I\subseteq[n]}\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq 2^{3m}$
-for any nonsignalling strategy pp. From 6.3, we obtain for $|I|>m$ that
+for any nonsignalling strategy $p$. From 6.3, we obtain for $|I|>m$ that
 $\sum_{\mathbf{x},\mathbf{y}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$ $=\sum_{\mathbf{y}}\sum_{\mathbf{x}|_{I^{c}}\in\{0,1\}^{|I|^{c}}}\sum_{\mathbf{x}|_{I}\in\{0,1\}^{|I|}}p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$
@@
 since there are $2^{2m}$ different $\mathbf{x}$ and $\mathbf{y}$, respectively. This finishes the proof. ∎
-6.3. Values for the 33-fold parallel repetition
-Now, we will look at the case of 33-fold parallel repetition.
+6.3. Values for the $3$-fold parallel repetition
+Now, we will look at the case of $3$-fold parallel repetition.
 We will start with describing a classical strategy with winning probability $\frac{5}{16}$.
-We note that this strategy performs better that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$.
+We note that this strategy performs better that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the $2$-fold parallel repetition, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$.
 Alice’s and Bob’s strategies are of the form $\left(a_{1},a_{2},a_{3}\right)$, $\left(b_{1},b_{2},b_{3}\right)$ where
@@
 $x_{1}=y_{2}\wedge y_{3},$ $y_{2}=x_{1}\vee x_{3}.$
-If $x_{1}=y_{2}\wedge y_{3}=0$, the players win round 33 if and only if $y_{2}=x_{3}$.
+If $x_{1}=y_{2}\wedge y_{3}=0$, the players win round $3$ if and only if $y_{2}=x_{3}$.
 But this is true because $y_{2}=x_{1}\vee x_{3}=x_{3}$ for $x_{1}=0$.
@@
 By exhaustive search through all deterministic strategies using a computer, we have confirmed that this strategy is indeed optimal.
-While we see that the previous classical strategy performs better than the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition, there is no non-signalling strategy that outperforms the one where we use optimal non-signalling strategy for one game and the optimal classical strategy for the 22-fold parallel repetition.
-Therefore, the non-signalling value of the 33-fold parallel repetition is $\frac{1}{3}$.
+While we see that the previous classical strategy performs better than the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the $2$-fold parallel repetition, there is no non-signalling strategy that outperforms the one where we use optimal non-signalling strategy for one game and the optimal classical strategy for the $2$-fold parallel repetition.
+Therefore, the non-signalling value of the $3$-fold parallel repetition is $\frac{1}{3}$.
 The following is an optimal strategy achieving this winning probability.